### v1

# Chenjie Zhong    Zhipeng Li†,*    Shangzhi Xu Xiaohu Li

Abstract
Variational Mode Decomposition (VMD) plays an important role in many scientific areas, especially for the area of signal processing. Unlike the traditional Fourier paradigm, it makes decomposition of a signal possible without any predefined function basis, which gives unprecedented flexibilities while handling narrow-banded signals of varieties. However, determining the number and central frequencies of intrinsic mode functions are still open questions in that few studies has been proposed to give a complete method that can theoretically guarantee the global convergence during decomposition. In this article, we propose a globally convergent numerical optimization method based on variational convex optimization to automatically determine the number and central frequency of IMFs without any prior knowledge for narrow banded signals, as long as the spectra of each sub-band are identifiable. Our method focuses on finding the support baseline of the spectral function, and further separating the significant frequency band regions above the support baseline. Unlike pioneer works that focus on optimizations on complex field, our method achieves obtaining the number of decomposed IMFs and center frequencies in real field by combining variational calculus, convex optimization, and numerical solutions of differential equations in real field and theoretical analysis shows our algorithm is guaranteed terminate to one of the optimum as close as possible. Experiments also shows that our algorithm converges quickly and can be used in practical engineering to determine the prior information of the number of intrinsic modes and evaluate initial center frequency as prior information to continue the VMD procedure.
keywords:
narrow-banded signals,variational optimization, Intrinsic Mode Function(IMF)
{MSCcodes}
65K10, 65L10, 68U01, 46N10, 34B05
1 Introduction
Variational Mode Decomposition (VMD) [5] decomposes original signals into multiple Intrinsic Mode Functions (IMF) with limited bandwidth with the help of variational optimization. The core of VMD is to minimize the sum of the bandwidths with respect to the estimated IMFs as well as to ensure the sum of all the IMFs is equal to the original signal. The paradigm of VMD is a significant progress in the field of signal processing in that it not only inherits the core of EMD that breaks the traditional Fourier-transform alike framework for signal decomposition where a signal must be decomposed in terms of fixed signal basis like sine and cosine functions, but is also considered as the first trial to solve the signal decomposition problem with the help of variational optimization. VMD plays an important role in modern signal decomposition after proposition. In the field of mechanics, [14] uses VMD to decompose the vibration signals of planetary gearboxes under variable-speed conditions, in biomedical field, VMD can be used to diagnose the seizure of epileptic by retrieving the feature of decomposed Electroencephalography (EEG) signals [21] and can also help to denoise the Electrocardiogram (ECG) signals [9]. In the field of Earth Physics signals, it not only helps to increase the Signal-to-Noise Ratio(SNR) by combining with adaptive non-local mean filtering [7], but also helps the prediction of monthly runoff by decomposing the signal into trend, periodic and random components as the input feature of downside neural networks [16]. Besides, VMD also provides the potential for the channel identification in communications [1] and helps to the visualization of the protein structure [15]. We show the detail of VMD in Supplementary Materials Section 1. Although VMD has provided flexibility in decomposing signals, there are factors that hinder its further applications in that the number of intrinsic mode function must be determined manually as a piror [19, 10, 18, 17] in coordination with penalty factor without a common principle. Besides, VMD procedure works in complex field and to our knowledge there is no rigorous analysis on the convergence[5]. In order to alleviate the problem, the subsequent research either needs the knowledge to the range of the number of modes [11], or retrieves the IMFs recursively [12], or highly dependent on a predefined window-size to retrieve the peak and valley of spectra [6]. To our best knowledge, although there are some research focusing on wide-band extension of VMD [22], yet even to narrow-band signals, there are almost no complete research that can simultaneously be rigorous in theoretical analysis, while obtaining estimation of the number of modes and center frequency without any prior knowledge to the signal, and achieving non-recursive iterations.
In this article, we propose a variation-optimization based, globally convergent method to automatically determine the IMFs numbers and the corresponding center frequencies for narrow-banded signals. In short, we model the modes retrieving problem as adaptively extracting the supporting baseline of the signal spectrum amplitude function. By subtracting the extracted baseline from the signal spectrum amplitude function, we retrieve the significant frequency band components of the signal and extract the various modes of the narrow-band signal. The extracted supporting baseline is acquired by adversarially maximizing the integral of the baseline while minimizing the integral of its curvature. The supporting baseline is computed by iteratively solving a variational problem that is degenerated into the solution of a normal ordinary differential equation, and under proper condition and weight parameters the method always converges to the optimal curve that can be used to determine the center frequencies and the number of IMFs after proper estimation of the gap residual at the bottom. Once the number of IMFs and initialized frequencies are determined by our method, these parameters can be used to start the following VMD procedure. Experiments show that our method can quickly get accurate estimation about the number of IMFs and the corresponding initial center frequencies as prior to the following VMD procedure.
Our contribution can be summarized as 3 folds. First, as far as we know, in determining the number and central frequencies of IMFs, we are the first one to not only give a feasible algorithm that can decompose a narrow-banded signal automatically but also prove the global convergence in a rigorous mathematical view. Second, our proof generalizes the classical dual ascent method into functionals, which expands the scope of the application for the classical algorithm. Third, experiments show that our algorithm can filter out all the potential modes without any prior setting or assumptions to the number of modes, and also show good performance in both precision and speed even when the center frequencies of different modes are very close.
This article will be organized as follows: In the following section we propose the main framework of our algorithm and give the proof of the convergence. In section 3 we introduce the implementation detail about the whole workflow. In Section 4 we design a series of experiments to evaluate the performance of our algorithm. In the final section we show our conclusion about the main contribution of the paper.
2 Methodology
In this section we introduce our method for estimating the number of IMF modes as well as center frequencies of each band. We first introduce the background of the problem and give a brief introduction about Variational Mode Decomposition, and then we propose the concept of Supporting Baseline of a Function, finally we describe our method for finding the baseline and the whole workflow.
2.1 Supporting Baseline of a Function
Without loss of generality we first introduce the Supporting Baseline of a Function. For any function $s(x)>0$, the supporting baseline is a curve that tightly fits the lower bound of the function and reflects the overall trend of the lower bound. Fig. 1 gives an intuitive illustration about the conception of supporting baseline. As illustrated in Fig. 1, intuitively the supporting baseline should be tightly close to the lower bound of the function, approaching the infimum of functions as close as possible as well as maintain its own smoothness in order to capture the global trend of the lower bound rather than too much details, as capturing too much details will in turn reduce the robustness for separating the modes since there might be some noise added to the signal in either time or frequency domain.
Figure 1: Illustration to the concept of supporting baseline: In (a)-(c) the solid line is $y=10\cos\pi x-10\cos 5\pi x+x^{2}$. The dashed line in (a) is a proper supporting baseline since it actually fits the lower bound of the solid line as well as ignoring high frequent details of the solid line. The dashed line in (b) has a gap to the bottom of the solid line so it is not suitable to be a good supporting baseline. In contrast, the dashed line in (c) fits too tightly to the lower bound of the solid line, capturing too much detail of the solid line so that it can be reckoned as the lower envelope rather than a proper supporting baseline.
2.2 Variational Methods for Finding the Supporting Baseline
To solve the problem for finding the supporting baseline, we use optimization method mainly based on variation and convex optimization [8, 4]. Let $f(x)\geq 0$ be a known function, the problem of finding the supporting baseline curve of $f(x)$ is equivalent to finding a function $g(x)$ that tightly approaching the infimum of $f(x)$ while maintaining the global trend of its lower bound. This can be viewed as the following optimization problem:
[…]
Converting (1) into standard form, we get:
[…]
This is an optimization problem with inequality constraints. To solve the problem, one can construct an Lagrange Slack Function by adding functional-styled weights to each factor and constraints
$L(g,\lambda,\mu,\alpha,\beta)$ $=\int_{\Omega}\alpha(x)g^{\prime\prime 2}(x)\text{d}x+\int_{\Omega}\beta(x)[-g(x)]\text{d}x$
(3) $+\int_{\Omega}\lambda(x)[g(x)-f(x)]\text{d}x+\int_{\Omega}\mu(x)[-g(x)]\text{d}x$
where
[…]
and $\alpha(x),\beta(x)>0$. The solution of the problem can be found by solving a variation problem, namely, if we set
(6) $F=\alpha(x)g^{\prime\prime 2}(x)-\beta(x)g(x)+\lambda(x)[g(x)-f(x)]-\mu(x)g(x)$
then the above Lagrange Slack Function is like the form $J[g(x)]=\int_{\Omega}F(x,g,g^{\prime\prime})\text{d}x$ so that if we assume the variation is zero at both ends of the curve, one can easily know the optimal solution of the above function happen to satisfy the simplified form of Euler-Poisson equation [8] if we let the variation vanish, so it can be written as
(7) $F_{g}+\cfrac{\text{d}^{2}F_{g^{\prime\prime}}}{\text{d}x^{2}}=0$
where
(8) $F_{g}=-\beta(x)+\lambda(x)-\mu(x)$
(9) $F_{g^{\prime\prime}}=2\alpha(x)g^{\prime\prime}(x)$
It is easily to get that
(10) $\cfrac{\text{d}^{2}F_{g^{\prime\prime}}}{\text{d}x^{2}}=2\alpha^{\prime\prime}(x)g^{\prime\prime}(x)+4\alpha^{\prime}(x)g^{(3)}(x)+2\alpha(x)g^{(4)}(x)$
By substituting (8) and (10) into (7), we can get the following ordinary differential equation
(11) $2\alpha^{\prime\prime}(x)g^{\prime\prime}(x)+4\alpha^{\prime}(x)g^{(3)}(x)+2\alpha(x)g^{(4)}(x)=\beta(x)-\lambda(x)+\mu(x)$
This is a 4th-ordered Ordinary Differential Equation (ODE), so that normally four constraints are needed to determine a definite solution. Recall that we assume the variation is zero at both ends of the curve, so that we can typically set four constraints at both ends of the curve with respect to the function value and its first-order derivative as follows:
(12) $g(0)=f(0),g(n)=f(n),g^{\prime}(0)=f^{\prime}(0),g^{\prime}(n)=f^{\prime}(n)$
where 0 stands for the left end and $n$ the right end of the curve. To solve the boundary value problem of ODE, we use Finite Difference Method [20]. Note that for a differential function $g(x)$, it’s derivative of first to fourth order can be expressed approximately as the linear combination of $g(x)$ itself and we can take similar operation to the function $\alpha(x)$ (See Supplementary Materials 2). Since now different order of derivative with respect to $\alpha(x)$ and $g(x)$ can be expressed as the linear combination of themselves, one can easily transform (11) into the matrix-form as follows:
(13) $(2\boldsymbol{A^{(2)}\alpha})\odot(\boldsymbol{G^{(2)}g})+(4\boldsymbol{A^{(1)}\alpha})\odot(\boldsymbol{G^{(3)}g})+(2\boldsymbol{\alpha})\odot(\boldsymbol{G^{(4)}g})=\boldsymbol{\beta-\lambda+\mu}$
where $\boldsymbol{A}^{(n)}$ and $\boldsymbol{G}^{(n)}$ is the conversion matrices (See Supplementary Materials 2) that transform the nth derivatives of $\alpha(x)$ and $g(x)$ into linear expression of the corresponding function, together with the boundary condition to form full-rank matrices, and $\boldsymbol{\alpha}$, $\boldsymbol{g}$, $\boldsymbol{\beta}$, $\boldsymbol{\lambda}$, $\boldsymbol{\mu}$ is obtained by respectively discretizing the corresponding functions, $\odot$ stands for component-wise multiplication between matrices(See Supplementary Materials 3 for details). The conversion matrices $\boldsymbol{A}^{(n)}$ and $\boldsymbol{G}^{(n)}$ share the same structure, in which the last four lines depicts the boundary condition and the other the ODE equation, and $\boldsymbol{g}$, $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$, $\boldsymbol{\lambda}$, $\boldsymbol{\mu}$ are column vectors in which each element is a uniformly discrete sampled value on the function $g(x)$, $\alpha(x)$, $\beta(x)$, $\lambda(x)$, $\mu(x)$, respectively.
2.3 Proof of the Convergence
Algorithm 1 The Projection Gradient Descent Method for Evaluating Support Baseline
1: Input $\{f(x)\}$
2: Initialize $\alpha(x),\beta(x),\lambda(x),\mu(x),g(x)$
3: Pre-process by extending $f(x)$ at both ends.
4:repeat
5:  $n\leftarrow n+1$
6:  Update $g(x)$:
7:   Solve the equation:
8:   $(2\boldsymbol{A^{(2)}\alpha})\odot(\boldsymbol{G^{(2)}g})+(4\boldsymbol{A^{(1)}\alpha})\odot(\boldsymbol{G^{(3)}g})+(2\boldsymbol{\alpha})\odot(\boldsymbol{G^{(4)}g})=\boldsymbol{\beta-\lambda+\mu}$
9:  Dual ascent:
10:   $\lambda^{n+1}\leftarrow\max\{0,\lambda^{n}+\theta(g-f)\}$
11:   $\mu^{n+1}\leftarrow\max\{0,\mu^{n}+\gamma(-g)\}$
12:until convergence: $\|g^{n}-g^{n-1}\|_{2}^{2}/\|g^{n}\|_{2}^{2}<\epsilon$
In this section we show the global convergence of the Algorithm 1. In our algorithm we have in fact solved a functional Extreme Value Problem with respect to the spectrum to the original signal. Despite the fact that the Fourier transform spectrum of a general signal may not necessarily be smooth, yet on one side, as in practical digital signal processing, the signal are discretized and can be smoothly interpolated, indicating the function processed by our algorithm are practically smooth which enable the reasonable use of commonly used assumptions that are advantageous for proving algorithm convergence, and on the other hand, subsequent numerical experiments have also shown that such assumptions not only do not affect the range of signals that can be processed, but are also very helpful for proving convergence. Therefore, in our proof, we assume that the spectrum is smooth so that we can extend the Lagrangian dual gradient ascent to functionals. Our proof combines functional, variational optimization and ODE theories in the framework of gradient projection, and can be divided into 3 steps. First we prove that we are dealing with a convex functional optimization problem, where the feasible domain is also a convex set. Next we show that under the condition of our algorithm, the Lagrangian dual gap is zero [4, 2] so that it is possible to find the optimum by dual ascent method with some modification from vectors to functional. After that, the convergence of iteration in dual space is shown and it is also shown that the optimal function can also be approached during the iterations by utilizing the dependence of differential equation solutions on coefficients. For the sake of limited spaces, we move the proof of convexity and duality into Supplementary Materials 4 and only give the proof to the convergence of iteration in Algorithm 1 in this part.
We first restate the iteration in Algorithm 1. Notice that during the iteration, we have
(14) $\left\{\begin{aligned} g_{k+1}(x)&=\underset{g(x)}{\arg\min}\mathcal{L}(g(x),\lambda_{k}(x),\mu_{k}(x))\\ \lambda_{k+1}(x)&=\max[0,\lambda_{k}(x)+\theta(g_{k+1}(x)-f(x))]\\ \mu_{k+1}(x)&=\max[0,\mu_{k}(x)+\gamma(-g_{k+1}(x))]\end{aligned}\right$.
to show the projection property for the $\max$ operation we first prove the following lemma:
Lemma 2.1.
Let $m(x),x\in\mathcal{D}$ be any continuous function that has both positive and negative values, Let $\Pi$ be an operator that makes all the negative values for $m(x)$ zero and non-negative unchanged, or $\Pi m(x)=\max(0,m(x))$, let $M(x),x\in\mathcal{D}$ be any non-negative function, then
(15) $\int_{\mathcal{D}}\bigl[m(x)-\Pi m(x)\bigr]\bigl[M(x)-\Pi m(x)\bigr]dx\leq 0$
Proof 2.2.
By expanding the equation (15) we should have the following:
(16) $\begin{split}&\ \ \ \ \int_{\mathcal{D}}\bigl[m(x)-\Pi m(x)\bigr]\bigl[M(x)-\Pi m(x)\bigr]dx\\ &=\int_{\mathcal{D}}m(x)M(x)dx-\int_{\mathcal{D}}\Pi m(x)M(x)dx-\int_{\mathcal{D}}m(x)\Pi m(x)dx+\int_{\mathcal{D}}\bigl[\Pi m(x)\bigr]^{2}dx\\ &=\int_{\mathcal{D}}m(x)M(x)dx-\int_{\mathcal{D}}\Pi m(x)M(x)dx-\int_{\mathcal{D}}\bigl[\Pi m(x)\bigr]^{2}dx+\int_{\mathcal{D}}\bigl[\Pi m(x)\bigr]^{2}dx\\ &\leq\int_{\mathcal{D}}m(x)M(x)dx-\int_{\mathcal{D}}m(x)M(x)dx-\int_{\mathcal{D}}\bigl[\Pi m(x)\bigr]^{2}dx+\int_{\mathcal{D}}\bigl[\Pi m(x)\bigr]^{2}dx=0\end{split}$
Moreover, if $m(x)$ is in feasible set so that $m(x)=\Pi m(x)$, then (16) will become zero.
By using Lemma 15, we can prove that in iteration (14) the multiplier $\lambda_{k}(x)$, $\mu_{k}(x)$ always terminate at one of the optimum. To say in detail, we should have the following theorem:
Theorem 2.3.
Let $\mathcal{G}(\lambda(x),\mu(x))=\underset{g(x)}{\arg\min}\mathcal{L}(g(x),\lambda(x),\mu(x))$ and $\lambda_{k}(x)$, $\mu_{k}(x)$ be equicontinuous for all $k=1,2,3,\cdots$, then if $\mathcal{G}$ satisfies
[…]
for any $\lambda_{1}(x)$, $\lambda_{2}(x)$, $\mu_{1}(x)$, $\mu_{2}(x)$, where $D$ stands for the variation, and the step size $\eta<\frac{1}{L}$, then as $k\rightarrow+\infty$, the iteration series of $(\lambda_{k}(x),\mu_{k}(x))$ must terminate at the point as close as possible to one of the optimum $(\lambda^{*}(x),\mu^{*}(x))$ for $\mathcal{G}$.
Proof 2.4.
For conciseness, simplifying the integral by introducing matrix-form expression, where
(21) $\left[\begin{array}[]{cc}a(x)&b(x)\end{array}\right]\left[\begin{array}[]{cc}c(x)\\ d(x)\end{array}\right]=\int_{\mathcal{D}}a(x)c(x)\text{d}x+\int_{\mathcal{D}}b(x)d(x)\text{d}x$
we have
(25) $\left[\begin{array}[]{cc}\tilde{\lambda}_{k+1}(x)-\lambda_{k+1}(x)&\tilde{\mu}_{k+1}(x)-\mu_{k+1}(x)\end{array}\right]\left[\begin{array}[]{cc}\lambda_{k}(x)-\lambda_{k+1}(x)\\ \mu_{k}(x)-\mu_{k+1}(x)\end{array}\right]\leq 0$
[…]
On the other side, for any functional we have
$\mathcal{G}(\lambda_{k+1}(x),\mu_{k+1}(x))=\mathcal{G}(\lambda_{k}(x),\mu_{k}(x))$
$+DG_{(\lambda_{k}(x),\mu_{k}(x))}(\lambda(x),\mu(x))^{\top}\begin{bmatrix}\lambda_{k+1}(x)-\lambda_{k}(x)\\ \mu_{k+1}(x)-\mu_{k}(x)\end{bmatrix}$
$+\frac{1}{2!}\begin{bmatrix}\lambda_{k+1}(x)-\lambda_{k}(x)&\mu_{k+1}(x)-\mu_{k}(x)\end{bmatrix}$
(31) $\quad D^{2}G_{(\lambda(x),\mu(x))}(\varepsilon_{k}(x),\zeta_{k}(x))\begin{bmatrix}\lambda_{k+1}(x)-\lambda_{k}(x)\\ \mu_{k+1}(x)-\mu_{k}(x)\end{bmatrix}$
where $(\varepsilon_{k}(x),\zeta_{k}(x))$ locates on general line segment from $(\lambda_{k}(x),\mu_{k}(x))$ to
$(\lambda_{k+1}(x),\mu_{k+1}(x))$. If $\mathcal{G}$ is second-order differentiable and satisfies (17) so that $D^{2}\mathcal{G}$ is bounded with [L,L][-L,L]. Due to the concavity of $\mathcal{G}$ and the left side of the bound, and combine (30) we get
[…]
if $\eta=\max\{\theta,\gamma\}=\cfrac{1}{L}<\cfrac{2}{L}$, then $\mathcal{G}$ is a monotonously increasing functional series with upper bound due to its concavity, so that $\mathcal{G}$ converges and the difference between adjacent term of both $\mathcal{G}$ and $(\lambda_{k}(x),\mu_{k}(x))$ is approaching zero.
Now we prove $\{g_{k}(x)\}$ is guaranteed to approach one of the optimum when the algorithm terminates even if $\mathcal{G}$ is not strongly concave. Due to the concavity the iteration on $\mathcal{G}$ should satisfy [2]
$l\left(\begin{bmatrix}\lambda_{k+1}(x)\\ \mu_{k+1}(x)\end{bmatrix};\begin{bmatrix}\lambda_{k}(x)\\ \mu_{k}(x)\end{bmatrix}\right)-\cfrac{1}{2\eta}\left\|\begin{bmatrix}\lambda_{k+1}(x)-\lambda_{k}(x)\\ \mu_{k+1}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
$\geq l\left(\begin{bmatrix}\lambda^{*}(x)\\ \mu^{*}(x)\end{bmatrix};\begin{bmatrix}\lambda_{k}(x)\\ \mu_{k}(x)\end{bmatrix}\right)-\cfrac{1}{2\eta}\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
(35) $+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}$
since $\mathcal{G}$ is convex. Note that the spectrum of $D^{2}\mathcal{G}$ is larger than L-L and $\eta=\cfrac{1}{L}$ to replace the left side of (35), and use the concavity of $\mathcal{G}$, we have
$\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}\leq 2\eta(\mathcal{G}(\lambda_{k+1}(x),\mu_{k+1}(x))$
(36) $-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x)))+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
Sum up (36) for each $k$ and eliminate the repeated terms that occur at the both ends of the inequality, we have
$\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}\leq 2\eta\sum\limits_{i=0}^{k}(\mathcal{G}(\lambda_{i+1}(x),\mu_{i+1}(x))$
(37) $-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x)))+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{0}(x)\\ \mu^{*}(x)-\mu_{0}(x)\end{bmatrix}\right\|_{2}^{2}$
As the left side is non-negative, and recall in (34) we know that $\{\mathcal{G}_{k}\}$ is non-decreasing and $\mathcal{G}^{*}$ is the supreme of $\mathcal{G}$, hence $\mathcal{G}(\lambda_{k+1}(x),\mu_{k+1}(x))-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x))$ is the largest among all the terms involved in the summation, by applying reasonable inequality scaling we must have
(38) $0\geq\mathcal{G}(\lambda_{k+1}(x)$ $,\mu_{k+1}(x))-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x))\geq-\cfrac{1}{2\eta(k+1)}\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{0}(x)\\ \mu^{*}(x)-\mu_{0}(x)\end{bmatrix}\right\|_{2}^{2}$
which shows the iteration is approaching the optimum. From (36) and by utilizing the monotonicity of the functional $\mathcal{G}$ during iteration we have
(39) $\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}<\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
which shows the distance between the iterating points and the optimum is always decreasing, since the distance is always non-negative so that it has lower bound, $(\lambda_{k}(x),\mu_{k}(x))-(\lambda^{*}(x),\mu^{*}(x))$ is a convergent sequence so that $(\lambda_{k}(x),\mu_{k}(x))$ is bounded. According to Arzelà-Ascoli Theorem, there exists some sub-sequences $(\lambda_{k_{i}}(x),\mu_{k_{i}}(x))$ that converges, so that we cannot assert that the sequence is necessarily convergent in all cases and need careful discussion. One one hand, if there is only one optimum in $\mathcal{G}$, then it is trivial to show that all these sub-sequences converges to the unique optimum by the constraints in (38). If $\mathcal{G}$ has more than one optimum, then suppose there exists more than one sub-sequence, for example, $(\lambda_{k_{i}}(x),\mu_{k_{i}}(x))$, $(\lambda_{k_{j}}(x),\mu_{k_{j}}(x))$ that converges to $(\lambda^{*}_{i}(x),\mu^{*}_{i}(x))$, $(\lambda^{*}_{j}(x),\mu^{*}_{j}(x))$ respectively, then due to (34) we show that although the sequence might have converged subsequence that approaches to different optimum, the adjacent term tends to zero so that the adjacent primal solution in iteration also tends to zero. On the other hand, as numerical algorithms will introduce error inevitably, if after some iteration $(\lambda_{k}(x),\mu_{k}(x))$ is exactly the unique optimum so that the right hand of (39) is zero, then from monotonicity of iteration we know $(\lambda_{k+1}(x),\mu_{k+1}(x))=(\lambda_{k}(x),\mu_{k}(x))$. Even if $\mathcal{G}$ has more than one optimum and after some iteration $(\lambda_{k}(x),\mu_{k}(x))$ approaches one of the optima $(\lambda^{*}_{1}(x),\mu^{*}_{1}(x))$, since there is another optimum $(\lambda^{*}_{2}(x),\mu^{*}_{2}(x))$, then we have
(40) $\mathcal{G}(m\lambda^{*}_{1}(x)$ $+(1-m)\lambda^{*}_{2}(x),m\mu^{*}_{1}(x)+(1-m)\mu^{*}_{2}(x))$
$=\mathcal{G}(\lambda^{*}_{1}(x),\mu^{*}_{1}(x))=\mathcal{G}(\lambda^{*}_{2}(x),\mu^{*}_{2}(x)),m\in[0,1]$
since the direction of the line connecting the median point and the initial point in the first-order Taylor expansion formula is consistent with $\left[\begin{array}[]{c}\lambda^{*}_{2}(x)-\lambda^{*}_{1}(x)\\ \mu^{*}_{2}(x)-\mu^{*}_{1}(x)\end{array}\right]$, so we have
(41) $D\mathcal{G}_{(\lambda_{k}(x),\mu_{k}(x))}(\lambda(x),\mu(x))\left[\begin{array}[]{c}\lambda^{*}_{2}(x)-\lambda^{*}_{1}(x)\\ \mu^{*}_{2}(x)-\mu^{*}_{1}(x)\end{array}\right]=0$
thus the variation direction is always orthogonal to any lines connected between different optima, as a consequence, when projected again during the next iteration, $(\lambda_{k+1}(x),\mu_{k+1}(x))$ must overlap with $(\lambda_{k}(x),\mu_{k}(x))$, so we have $(\lambda_{k+1}(x),\mu_{k+1}(x))=(\lambda_{k}(x),\mu_{k}(x))$. In all cases we show $(\lambda_{k}(x),\mu_{k}(x))$ is always convergent to at least one of the optimum and stay still as long as it arrives the optimum during the iteration even if there are more than one optimum in $\mathcal{G}$.
Now we turn to prove the convergence of series $g_{k}(x)$. As each $g_{k}(x)$ is obtained by solving an ODE function as the optimum with respect to a variational problem, we should first prove the existence and uniqueness to the solution of $g$ for each iteration.
Theorem 2.5.
The solution of (11) with $g(0)=0,g(n)=0,g^{\prime}(0)=0,g^{\prime}(n)=0$ is $g(x)\equiv 0$.
Proof 2.6.
Note that the differential equation of the boundary value problem we are solving has only zero solutions under the corresponding homogeneous differential equation and homogeneous boundary conditions, in order to prove the existence and uniqueness of our solution in each iteration [3]. Recall that when dealing with the homogeneous condition, we exactly have
(42) $g(x)=\int_{a}^{x}\left(\int_{a}^{x}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t\right)\text{d}t+C_{3}x+C_{4}$
where
(43) $g(a)=g(b)=g^{\prime}(a)=g^{\prime}(b)=0$
It is obvious that $C_{3}$ and $C_{4}$ is zero by $g(a)$ and $g^{\prime}(a)$, and $g^{\prime}(b)=0$ is equivalent to
(44) $\int_{a}^{b}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t=0$
If $\alpha(t)>0$ and $\alpha(t)\in C[a,b]$, since $C_{1}t+C_{2}$ is a linear function so that if the zero point is not in $[a,b]$ then it is not possible to make (44) hold. If there exists $t\in[a,b]$ that makes $C_{1}t+C_{2}$ zero, then the upper limit integral function $\Phi(x)=\displaystyle\int_{a}^{x}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t$ has no zero points in $(a,b)$, thus makes $g(b)=\displaystyle\int_{a}^{b}\Phi(t)\text{d}t$ non-zero, which is paradoxical to our assumption if not both $C_{1}$ and $C_{2}$ is zero. As a consequence, during each iteration, there always exist a unique solution to $g(x)$, so that we can simply discuss the convergence based on the continuous-dependency between the solution and the coefficients.
Corollary 2.7.
For each iteration of (13), there always exists a unique solution.
Now we begin to prove the convergence of $g_{k}(x)$. For this property during the iteration in Algorithm 1, we have the following theorem:
Theorem 2.8.
Let $f(x)$ be first-ordered Lipschitz-continuous function that satisfies $|f^{\prime}(x)-f^{\prime}(y)|\leq L|x-y|$ for any $x,y$, and suppose $g_{0}(x)$ satisfies $0\leq g_{0}(x)\leq f(x)$. If $\{g_{k}(x)\}$ satisfies
[…]
Then if the step $\eta$ is sufficiently small so that $\eta<\frac{1}{L}$ for each iteration, $g_{k}(x)$ must terminate at one of the optimum.
Proof 2.9.
Based on lemma 15, theorem 2.3 and theorem 2.5 we have shown that our original problem has unique solution during each iteration. As (11) is a linear ODE, its solution is continuously dependent on its coefficients, so we have $\lim_{k\rightarrow+\infty}g_{k}(x)=g^{*}(x)$ when $\{(\lambda_{k}(x),\mu_{k}(x))\}$ converges or falls early in one of the optima, and
[…]
$\{(\lambda_{k}(x),\mu_{k}(x))\}$ has sub-sequences approaching different optimum respectively. In all cases we can terminate algorithm and guarantee the $\{g_{k}(x)\}$ terminates to one of the optimum as close as possible.
3 Implementation
3.1 Algorithm
In this part we state the whole implementation of our method. Our implementation can be mainly divided into three parts. Given a signal to be decomposed, we first use variational method in terms of iteratively constructing such a function sequence that each term of it is the optimal solution obtained by vanishing the variation of the Lagrangian Slack Function under the current evolutionary state with respect to inequality constraint conditions and then updating the weight of constraint using dual ascent method. Thirdly, we apply Kernel Density Estimation (KDE) to the amplitude values of each point in the spectrum to eliminate residual gaps at the bottom, which is based on the observation that the residual gap at the bottom will results in a clustering behaviour around the corresponding value that the residual gap share. The proper threshold to eliminate the gap for further evaluating the number of IMFs should locate where the value of the KDE function achieves the maximum. In short, our whole workflow can be summarized as in Algorithm 2.
Algorithm 2 The Whole Workflow
1: Input $\{f(x)\}$
2: Initialize $\alpha(x),\beta(x),\lambda(x),\mu(x),g(x)$
3: Pre-process by extending $f(x)$ at both ends.
4:repeat
5:  Run the kernel process of Algorithm 1
6:until convergence: $\|g^{n}-g^{n-1}\|_{2}^{2}/\|g^{n}\|_{2}^{2}<\epsilon$
7: Estimate KDE for the computed $g(x)$.
8: Compute best threshold $t_{best}=\arg\max\text{KDE}_{g(x)}(t)$.
9: Subtract noise-gap threshold from $g(x)$ by $g(x)-t_{\text{best}}$.
10: Get $k$, $\{\omega_{k}\}$ from non-zero intervals.
11: Begin VMD procedure [5] with $k$ and $\{\omega_{k}\}$
12: Output $\{u_{k}\}$
3.2 Pre-processing
In order to solve the equation effectively, before we start the iteration, we first normalize the signal into $[0,1]$, in both domain and range. Since in each iteration, our optimal baseline should satisfy the boundary condition of the ODE, so that when dealing with high-pass or low-pass signals where spectrum has sharp slope at the ends, the to-be-evaluated baseline will also exhibits the sharp slope, which will in turn have bad effect on the convergence at the ends. To overcome the boundary effect, we extend our original function by making a gentle decay at both ends of the spectrum that not only completes the peaks at both ends but also alleviate the sharp slope effect. We will discuss this in detail later in our experiments.
3.3 Post-processing
Despite the fact that the iteration procedure to find the supporting baseline converges in reasonable time in most cases when applied with proper extension to the original spectrum of the signal, the shape of the spectrum do have great impact on the convergence rate. For example, since our optimization strategy discourage the curvature (via second-order derivative) of the curve, the optimal baseline is prone to converge more faster at points where original spectrum is flat. On the other side, although $g(x)$ will eventually adhere closely to the lower bound of $f(x)$, it is difficult to guarantee that the error between $f(x)$ and $g(x)$ calculated after stopping the iteration is consistent at every point so that f(x)g(x)f(x)-g(x) will have some non-zero points where it should be zero, which makes further retrieving modes of difficulty. Since $g(x)$ converges to the bottom of $f(x)$, it is reasonable to assume that the gap is typically small so that f(x)g(x)f(x)-g(x) is almost zero except for those frequency points corresponds with effective signal components. In order to eliminate the gap, we use Kernel Density Estimation(KDE) method to estimate the best threshold, so that the gap of f(x)g(x)f(x)-g(x) will be exposed where the value of the KDE function achieves the maximum. Fig. 2 illustrates such a phenomenon.
Figure 2: Illustration of the non-consistence of the error between the converged optimal supporting baseline and the original function. Fig. 2a shows the converged $g(x)$ and Fig. 2b shows f(x)g(x)f(x)-g(x). Note that since $g(x)$ does have gap to the lower bound of $f(x)$ so that the bottom of f(x)g(x)f(x)-g(x) is almost not on a horizontal line everywhere, which indicates there are some non-zero gaps. These non-zero gaps will have a significant impact on the stability of the subsequent modal number solution. 2c shows the estimate kernel density pertaining the gap at the bottom. 2d shows the estimated f(x)g(x)f(x)-g(x) after removal of the gap at the bottom.
4 Numerical Results
In this section we first show our experiments to our method. In order to prove the effectiveness of our method, we choose some standard function that were repeatedly introduced as benchmarks in some pioneer works [5, 13, 12], yet pay more attention on investigating the evolution of our algorithm during the course of iteration, to show the convergence and robustness of our method. Since our algorithm can be reckoned as an algorithm for finding the number of intrinsic modes and initial center frequencies that the proceeding VMD module needs, our experimental focus is on exploring the behaviour of the algorithm towards various signals in the process of finding the baseline and talking about the KDE process for eliminating the gap, rather than elaborating the performance brought by the subsequent VMD decomposition in the final decomposition results. Then we turn to discuss why we should and how we eliminate the boundary effect by extending the spectrum obtained from the FFT of original signals.
4.1 Experiments on Finding the Supporting Baseline
In this part we show our experiments on finding the supporting baseline for some benchmark signals that were frequently used to test the performance of algorithms in several articles [5, 12]. As default, all signals are obtained by linearly normalizing the domain of the original function to between 0 and 1, and the range of which is normalized into 0 and 1 before iterations to keep the stability. We uniformly add 200 points to the domain and only preserve half of the transformed spectrum as our initial spectrum to be iterated according to the properties of Fourier Transform with respect to real signals. In this way, the Nyquist frequency is 100Hz, which means the highest signal angular frequency that can be effectively represented is $200\pi$, which corresponds to 1 in our normalized FFT results. Our convergence criterion is $\cfrac{\|g_{n}-g_{n-1}\|}{\|g_{n}\|}<10^{-4}$.
We choose five typical signals including single mode signal, multi-mode signal, piecewise continuous signal, signal that are not necessarily narrow-banded, composite signal in which the center frequency of its components are very close. Due to the space limitations, we only illustrate the final convergent state of the iteration for the evaluation of the support baseline, as well as the convergent curve with respect to the error, and indicate the number of iteration steps in the caption. The intermediate iteration process status and the decomposed signal is given in detail in the supplementary materials 5.
Experiment 1: The first example is a single mode signal, on the other words, the signal is narrow-banded and only contains one center frequency. Here we choose $y(t)=100\sin 20\pi t$. As we see the normalized center frequency is $0.1$.
Figure 3: The initial, final convergent state and convergence curve in logarithmic coordinate system with respect to the error for finding the optimum supporting baseline of FFT with respect to time series $y(t)=100\sin 20\pi t$. The iteration lasts for 5161 steps and costs 1.88 seconds.
Experiment 2: The second example is a multi-mode signal, $y(t)=10\cos(10\pi t)+20\sin(20\pi t)$, which has two modes, exactly. The normalized center frequency is 0.05 and 0.1, which is relatively very close to each other. However, our algorithm can still survive where center frequencies of different modes are very close, since from 6000th iteration and on, the support baseline forms a small bulge at the close frequency points, which just separates the two frequency points.
Figure 4: The final convergent state in logarithmic coordinate system for finding the supporting baseline of FFT with respect to FFT of $y(t)=10\cos 10\pi t+20\sin 20\pi t$. The iteration lasts for 5258 steps and costs 1.89 seconds.
Experiment 3: The next example is a little complicated, it has four modes totally, we choose
(45) $\footnotesize y(t)=\left\{\begin{aligned} 6t^{2}+\cos(10\pi t+10\pi t^{2})+\cos(60\pi t),&\qquad t\in[0,0.5]\\ 6t^{2}+\cos(10\pi t+10\pi t^{2})+\cos(80\pi t-10\pi),&\qquad t\in(0.5,1]\end{aligned}\right$.
as our signal. It has a low frequency component, $6t^{2}$, two pure harmonic components, $\cos 60\pi t,\cos(80\pi t-10\pi)$, in different intervals that does not intersect with each other, respectively, and one narrow banded component, $\cos(10\pi t+10\pi t^{2})$, in the whole interval. To our surprise, the iteration converges even more quickly than those signals that has fewer modes. In our cases, it takes about 15000 iterations to achieve the convergence.
Figure 5: The final convergent state in logarithmic coordinate system for finding the supporting baseline of FFT with respect to time series (45). Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed supporting baseline precisely captured the trend of the lower bound so that all the modes can be filtered out. The iteration lasts for 3026 steps and costs 1.07 seconds.
Experiment 4: Now we show some examples on signals that conflict with the narrow-band assumption. We take $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$. As we see that, the spectrum of the signal exhibits two main peaks, however the signal is not a completely narrow-band signal since it exhibits comb-shaped spectrum with respect to harmonic components, although the energy of those harmonic components are relatively very small(Fig. 6). Our algorithm extracts three components of the signal, despite that in the mainstream view there should be only two IMFs. However, since it is not a narrow-band signal, the difficulty of decomposition objectively exists since one can not even provide a specific standard to determine whether these small harmonic components should be considered as an intrinsic mode. We also observe that, even in [5, 12], when carefully zoomed in, there does exist small oscillations in the first decomposed component which indicates the non-purity of it. Despite the fact that our results has three components, every separate component exhibits better purity than the pioneer works. On the other side, from the perspective of reconstruction error, it is shown in this example that, results generated by algorithm in [5] with both 2 and 3 modes has reconstruction error as twice as large as ours (0.1 ours vs 0.2 theirs, see S5.13a), which also indicates that our algorithm has better performance in this case.
Figure 6: The final convergent state in logarithmic coordinate system for finding the baseline of function $\frac{1}{1.2+\cos 2\pi t}+\frac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$. The iteration lasts for 3463 steps and costs 1.25 seconds.
Experiment 5: Next we show even when the central frequencies are very close, our algorithm can also work well. We take signals as $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$. The signal has one low-frequency component and 10 alternating current components with different amplitudes, of which the minimum difference between frequency gaps is only 0.04 after normalization and our algorithm can still survive in figuring out all the components.
Figure 7: The final convergent state in logarithmic coordinate system for finding the baseline of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The iteration lasts for 1768 steps and costs 0.64 seconds.
4.2 Experiments on Extension of Original Spectrum at Both Ends
Although supporting baseline can be retrieved by the method stated above, in real cases there are circumstances where convergence is very slow. For example, since the baseline must satisfy the boundary condition on both side, the trend of spectrum at both ends has a great impact to the shape of the iterating curve. In particular, for low or high pass signals of which the spectrum have a significant slope trend in low and high frequency regions, the iterating curve is prone to adhere the slope at both ends and continue to keep such inertia to some extent even when slipping out of those slope regions. This phenomenon brings us two bad effects in analysing signals: On one side, for low or high pass signals, their spectrum at both ends always exhibit half peaks since the scope of representable spectrum from FFT is always limited and the spectrum of those signals will inevitably hits the boundary, making the peak not complete and the estimation of center frequency inaccurate. On the other side, inertia effect will be harmful to the shape of the baseline, since in regions adjacent to low and high frequency bands, the function value of supporting baseline will have very large gap to the lower bound of the original spectrum due to the inertia of the downward trend. In order to circumvent such effect, we extend the original spectrum by forcing the spectrum to gently drop to the minimum of original spectrum and leave some extra flat intervals at both ends. This method not only take the advantage of trend effect, in that the curve will tend to keep horizontally at both ends which can make the peak be finally filtered out, but also complete the peaks at both ends without any harmful effects to the other regions as the experiment shows. Fig. 8 shows the evolution of spectrum baseline during the iteration with respect to the signal $y(t)=60t+60\cos 200\pi t$. This typical signal is a band-stop signal which shares both high and low pass feature. Note that if we don’t extend the spectrum, the baseline exhibits a sharp decreasing in the low frequency area and can not catch up the lower bound in the interval $[0.05,0.4]$ and the similar phenomenon also happens at the high-frequency area, even after 1500 iterations. This shows the necessity of proper extension to the original spectrum.
Figure 8: Iterations for finding the supporting baseline without extending the spectrum to the signal $y(t)=60t+60\cos 200\pi t$. The 500th(a), 1500th(b) iterations with no extension and the 500th(c), 1500th(d) iterations with extension are shown respectively. The solid line are the original spectrum and the dotted line are the evaluated supporting baseline during the iterations.
4.3 Experiments on Central Frequencies
In order to prove that our algorithm can find frequencies close to the practical central frequencies for each IMF component, we compare our results from the above Example 1 to Example 5, with the results of the original VMD process with the same number of IMF, to keep away the impact to the performance of VMD from the initial frequency value that we provide to the following VMD procedure. We first get the frequency estimation from our algorithm and the result of pure VMD procedure with the same number of IMFs respectively, and then sort them in order, to observe the difference between the results of ours and the corresponding values from VMD.
Table 1: The evaluated center frequencies
Experiment Ours Original VMD Max Difference
1 0.0550 0.0496 0.0054
2 0.0300, 0.0575 0.0251, 0.0507 0.0068
3 0.0100, 0.0775 0.1550, 0.2075 0.0008, 0.0530 0.1510, 0.2 0.0245
4 0.0125, 0.0875 0.2475 6.46e-5, 0.0067 0.0803 0.1682
5 0.005,0.0775 0.1, 0.125 0.1550, 0.1800 0.2075, 0.2325 0.2575, 0.2825 0.3075 0.000845, 0.0747 0.0999, 0.1250 0.1510, 0.182 0.223, 0.25 0.2760, 0.303 0.339 0.0315
Table 1 shows the result, where in each experiment the evaluated center frequencies for both ours and the pure VMD are sorted, and the max difference is computed as $\max\limits_{i}|f_{i}-F_{i}|$, where $f$ stands for our estimated frequencies and $F$ pure VMD’s. As shown in Table 1, in most cases, our estimation to the center frequencies has low difference with respect to the pure VMD. It can be found that, in Experiment 4, the error between estimated center frequencies of ours and pure VMD’s are larger than the average, however, as analysed in Experiment 4, on one side, the original VMD, no matter given IMFs number 2 or 3, fails to achieve reconstruction error as small as ours, on the other side, as shown in Fig.S5.13a(c) even when both started at IMFs number 3, the original VMD seems produce replicated modes at the Gaussian-alike component, with almost the same amplitude and shape, while ours provide relatively independent components when given our initial estimation for the center frequencies, which seems to indicates that our estimation in this case is more reliable.
5 Conclusion
In this article, we proposed a globally convergent method that can automatically determine the number of intrinsic mode functions for Variational Mode Decomposition. For narrow-banded signals, this method does not need any prior knowledge about the scope of the IMFs number, and provide non-recursively schema to determine IMFs and corresponding central frequencies simultaneously. Experiments show that, our method can provide accurate number of IMFs and corresponding center frequencies even if the center frequencies are very close. Our implementation is performant in both accuracy and speed. Moreover, rigorous proof is given for the convergence of the method, making it theoretically reliable.
6 Acknowledgement
This study is supported by the National Natural Science Foundation of China under Grant No. 61773290 and the Fundamental Research Funds for the Central Universities (22120230311).
SUPPLEMENTARY MATERIAL
1 Introduction of VMD
The key idea of VMD is to decompose an input signal into a series number of sub-signals via variational optimization method in spectral domain[5]. It is assumed the frequency of each mode is almost compact around a center pulsation and the associated analytic signal is computed by means of Hilbert Transform to get the unilateral frequency spectrum. Then each mode’s frequency spectrum was shifted into baseband by multiplying a frequency modulation factor with respect to the estimated center frequency, transforming each band-pass mode into low pass, and the bandwidth can be estimated by $H^{1}$ Gaussian smoothness. Afterwards, they decompose the original signal into each modes by minimizing the sum of the bandwidth of each mode with the constraints that the summation of all modes is the original signal. In short, their key variation optimization can be listed as following:
$\min\limits_{u_{k},\omega_{k}}\left\{\sum\limits_{k}\left\|\partial_{t}\left[\left(\delta(t)+\frac{\text{j}}{\pi t}\right)*u_{k}(t)\right]e^{-\text{j}\omega_{k}t}\right\|_{2}^{2}\right\}$
$\text{s.t.}\quad\sum\limits_{k}u_{k}=f$
so that each mode can be altered into an associated analytical signal by using Hilbert Transform and shifted into base-band for the following estimation of bandwidth using H1 Gaussian smoothness with the shifted demodulated components. To solve the problem, one can construct an augmented Lagrangian Function by introducing both quadratic penalty and Lagrangian Multipliers as follows:
$\mathcal{L}(u_{k},\omega_{k},\lambda)$ $=\alpha\sum_{k}\left\|\partial_{t}\left[\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)\right]e^{-j\omega_{k}t}\right\|_{2}^{2}$
$\quad\quad\quad\quad\quad\quad\quad+\left\|f-\sum u_{k}\right\|_{2}^{2}+\langle\lambda,f-\sum u_{k}\rangle$
This problem can be solved by using ADMM(Alternate Direction Method of Multipliers), in which each component can be updated by solving the equivalent minimization problem:
$u_{k}^{n+1}=\underset{u_{k}\in\mathbb{R}}{\arg\min}\left\{\alpha\left\|\partial_{t}\left[\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)\right]e^{-j\omega_{k}t}\right\|_{2}^{2}+\left\|f-\sum u_{i}+\frac{\lambda}{2}\right\|_{2}^{2}\right\}.$
One can rewrite it in frequency domain form, so that convolution becomes multiplication for the convenience of the calculation in subsequent variational operations:
$\hat{u}_{k}^{n+1}=\underset{\hat{u}_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\alpha\left\|j\omega\left[1+\text{sgn}(\omega+\omega_{k})\right]\hat{u}_{k}(\omega+\omega_{k})\right\|_{2}^{2}+\left\|\hat{f}-\sum\hat{u}_{i}+\frac{\lambda}{2}\right\|_{2}^{2}\right\}.$
and by substitution of $\omega\rightarrow\omega+\omega_{k}$ to perform a translation and considering the Hermitian symmetry of spectrum with respect to the real signals, both terms can be transformed into half-space integrals with non-negative frequencies:
$\hat{u}_{k}^{n+1}=\underset{u_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\int_{0}^{\infty}4\alpha(\omega-\omega_{k})^{2}|\hat{u}_{k}(\omega)|^{2}d\omega+2\int\left(\hat{f}-\sum\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)^{2}d\omega\right\}$
which can be evaluated by vanishing the first variation of positive frequencies for both signal components and center frequencies
$\hat{u}_{k}^{n+1}=\left(\hat{f}-\sum_{i\neq k}\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)\frac{1}{1+2\alpha(\omega-\omega_{k})^{2}}$
$\omega_{k}^{n+1}=\frac{\displaystyle\int_{0}^{\infty}\omega\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}{\displaystyle\int_{0}^{\infty}\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}$
2 Finite Discrete Expression and Their Matrices
Here we show finite discrete expression and the corresponding the conversion matrices with respect to (11),(12). Since the finite difference expressions for $\alpha(x)$ and $g(x)$ are identical in form, the conversion matrices $\boldsymbol{G}^{(n)}$ and $\boldsymbol{A}^{(n)}$ also share the exactly the same structure
(S2.1) $g^{(1)}(x)\approx\cfrac{-g(x+2h)+8g(x+h)-8g(x-h)+g(x-2h)}{12h}$
(S2.2) $g^{(2)}(x)\approx\cfrac{-g(x+2h)+16g(x+h)-30g(x)+16g(x-h)-g(x-2h)}{12h^{2}}$
(S2.3) $g^{(3)}(x)\approx\cfrac{g(x+2h)-2g(x+h)+2g(x-h)-g(x-2h)}{2h^{3}}$
(S2.4) $g^{(4)}(x)\approx\cfrac{g(x+2h)-4g(x+h)+6g(x)-4g(x-h)+g(x-2h)}{h^{4}}$
(S2.5) $\boldsymbol{G}^{(1)}=\cfrac{1}{12h}\begin{bmatrix}1&-8&0&8&-1&0&\cdots&\cdots&\cdots&0\\ 0&1&-8&0&8&-1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&1&-8&0&8&-1&0\\ 0&0&\cdots&\cdots&0&1&-8&0&8&-1\\ 12h&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&12h\\ -12&12&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-12&12\end{bmatrix}$
(S2.6) $\boldsymbol{G}^{(2)}=\cfrac{1}{12h^{2}}\begin{bmatrix}-1&16&-30&16&-1&0&\cdots&\cdots&\cdots&0\\ 0&-1&16&-30&16&-1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&-1&16&-30&16&-1&0\\ 0&0&\cdots&\cdots&0&-1&16&-30&16&-1\\ 12h^{2}&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&12h^{2}\\ -12h&12h&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-12h&12h\end{bmatrix}$
(S2.7) $\boldsymbol{G}^{(3)}=\cfrac{1}{2h^{3}}\begin{bmatrix}-1&2&0&-2&1&0&\cdots&\cdots&\cdots&0\\ 0&-1&2&0&-2&1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&-1&2&0&-2&1&0\\ 0&0&\cdots&\cdots&0&-1&2&0&-2&1\\ 2h^{3}&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&2h^{3}\\ -2h^{2}&2h^{2}&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-2h^{2}&2h^{2}\end{bmatrix}$
(S2.8) $\boldsymbol{G}^{(4)}=\cfrac{1}{h^{4}}\begin{bmatrix}1&-4&6&-4&1&0&\cdots&\cdots&\cdots&0\\ 0&1&-4&6&-4&1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&1&-4&6&-4&1&0\\ 0&0&\cdots&\cdots&0&1&-4&6&-4&1\\ h^{4}&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&h^{4}\\ -h^{3}&h^{3}&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-h^{3}&h^{3}\end{bmatrix}$
where the last four lines in matrices correspond to the boundary condition and the conversion matrices $\boldsymbol{A}^{(n)}$ have the same structure.
3 About Component-Wise Multiplication
Our component-wise multiplication is not necessarily the Hadamard production of matrices, in that we add some special rules in dealing with column vector, to say in detail, $\forall\boldsymbol{A}_{m\times n},\boldsymbol{v}_{m\times 1}$, we have
(S3.1) $\boldsymbol{v}\odot\boldsymbol{A}=\boldsymbol{A}\odot\boldsymbol{v}=\boldsymbol{A}*\boldsymbol{V}_{m\times n}$
where * stands for standard Hadamard production and each column of $\boldsymbol{V}$ is just a copy of $\boldsymbol{v}$. In this way, we can transform (13) into
(S3.2) $((2\boldsymbol{A^{(2)}\alpha})\odot\boldsymbol{G^{(2)}})g+((4\boldsymbol{A^{(1)}\alpha})\odot\boldsymbol{G^{(3)}})g+((2\boldsymbol{\alpha})\odot\boldsymbol{G^{(4)}})g=\boldsymbol{\beta-\lambda+\mu}$
By applying this rule we have $(\boldsymbol{A}\boldsymbol{a})\odot(\boldsymbol{B}\boldsymbol{b})=((\boldsymbol{A}\boldsymbol{a})\odot\boldsymbol{B})\boldsymbol{b}$ so that we can treat (13) as common matrix equation and get the solution.
4 Supplementary Material of the Proof
4.1 Proof of Convexity
Let $\mathcal{L}(g,\lambda,\mu,\alpha,\beta)=J_{1}(g(x))+J_{2}(g(x))$, where $J_{1}(g(x))=\displaystyle\int_{\Omega}\alpha(x)g^{\prime\prime 2}(x)\text{d}x$ and $J_{2}(g(x))=\displaystyle-\int_{\Omega}\beta(x)g(x)\text{d}x$. For $J_{1}(g(x))$, we have
(S4.1) $\lambda J_{1}(g_{1}(x))+(1-\lambda)J_{1}(g_{2}(x))-J_{1}[\lambda g_{1}(x)+(1-\lambda)g_{2}(x)]$
$=\lambda(1-\lambda)\int_{\Omega}\alpha(x)[g_{1}^{\prime\prime}(x)-g_{2}^{\prime\prime}(x)]^{2}\text{d}x\geq 0$
and $J_{2}(g(x))$ is obviously a linear funtional, so we have
(S4.2) $J[\lambda g_{1}(x)+(1-\lambda)g_{2}(x)]\leq\lambda J[g_{1}(x)]+(1-\lambda)J[g_{2}(x)]$
so that our objective functional is convex.
4.2 Proof of the Zero Duality Gap
Now we begin to prove that the duality gap is zero. Let
(S4.3) $\mathcal{O}(g(x))=\displaystyle\int_{\Omega}\alpha(x)g^{\prime\prime 2}(x)\text{d}x-\int_{\Omega}\beta(x)g(x)\text{d}x$
be the objective functional, and construct Lagrangian functional
$\mathcal{L}(g(x),\lambda(x),\mu(x))=\displaystyle\int_{\Omega}\alpha(x)g^{\prime\prime 2}(x)\text{d}x$ $-\int_{\Omega}\beta(x)g(x)\text{d}x$
(S4.4) $+\int_{\Omega}\lambda(x)[g(x)-f(x)]\text{d}x-\int_{\Omega}\mu(x)g(x)\text{d}x$
define dual functional as
(S4.5) $\mathcal{G}(\lambda(x),\mu(x))=\inf\limits_{g(x)\in\mathcal{D}}\mathcal{L}(g(x),\lambda(x),\mu(x))$
then on one side, we must have
(S4.6) $\mathcal{G}(\lambda^{*}(x),\mu^{*}(x)=\inf\limits_{g(x)}\mathcal{L}(g(x),\lambda^{*}(x),\mu^{*}(x))\leq\inf\limits_{g(x)}\mathcal{O}(g(x))=\mathcal{O}(g^{*}(x))$
where * stands for the optimum. On the other hand, one can construct general sets
(S4.7) $\mathcal{S}=\{(p(x),q(x),o)|g(x)-f(x)\leq p(x),-g(x)\leq q(x),\mathcal{O}(g(x))\leq o\}$
which forms an epi-graph to (2). thus
(S4.8) $\forall\xi(x)=\lambda(p_{1}(x),q_{1}(x),o_{1})+(1-\lambda)(p_{2}(x),q_{2}(x),o_{2})$
one can always find
(S4.9) $\eta(x)=\lambda g_{1}(x)+(1-\lambda)g_{2}(x)$
that makes
[…]
hold and thus $\mathcal{S}$ is a convex set. It is obvious that $(0,0,\mathcal{O}(g^{*}(x)))\in\partial S$, so according to the Support Theorem of Convex Sets there must be a general hyperplane that on one side pass through $(0,0,\mathcal{O}(g^{*}(x)))\in\partial S$ and on the other side ensures any point in $\mathcal{S}$ locates above the hyperplane. That is, there exists $\lambda(x),\mu(x)$ that makes
(S4.11) $\int_{\Omega}\lambda(x)p(x)\text{d}x+\int_{\Omega}\mu(x)q(x)\text{d}x\geq\mathcal{O}(g^{*}(x))-o$
hold for all $(p(x),q(x),o)$ in $\mathcal{S}$. Note that at least $\forall p(x),q(x)\geq 0$, $\mathcal{O}(g^{*}(x))\geq o$, so that the right end in this case is always non-negative. To summarize all above, we must have $\lambda(x)\geq 0$ and $\mu(x)\geq 0$, which indicates we can find a solution in dual space making the dual-gap vanish.
Next we prove the convexity of our feasible domain. The feasible set is $0\leq g(x)\leq f(x)$, so that $\forall g_{1}(x),g_{2}(x)\in[0,f(x)]$, we have $\lambda g_{1}(x)+(1-\lambda)g_{2}(x)\in[0,f(x)],\forall\lambda\in[0,1]$, which means the feasible set is also convex.
5 Experiment Detail
In this section we begin to provide some details of our experiments as complementary.
Figure S5.1: The iteration for finding the supporting baseline of FFT with respect to time series $y(t)=100\sin 20\pi t$. The 1000th(a), 2000th(b), 3000th(c), 4000th(d), 5000th(e) iteration are respectively shown.
Figure S5.2: The gap-elimination course for the found spectrum supporting line with series $y(t)=100\sin 20\pi t$. (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
Figure S5.3: The decomposition result of $y(t)=100\sin 20\pi t$ from the subsequent VMD, using the found number of modes and center frequency as initialization. The subfigure from top to bottom are respective the source signal, the decomposed mode(only one) and the residual.
Figure S5.4: The iteration for finding the supporting baseline of FFT with respect to time series $y(t)=10\cos(10\pi t)+20\sin(20\pi t)$. The 1000th(a), 2000th(b), 3000th(c), 4000th(d), 5000th(e) iterations are respectively shown.
Figure S5.5: The gap-elimination course for the found spectrum supporting line with series $y(t)=10\cos 10\pi t+20\sin 20\pi t$. (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
Figure S5.6: The decomposition result of $y(t)=10\cos 10\pi t+20\sin 20\pi t$ from the subsequent VMD, using the found number of modes and center frequency as initialization. The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are two) and the residual.
Figure S5.7: The iteration for finding the supporting baseline of FFT with respect to time series (45). The 1000th(a), 2000th(b), 3000th(c) iterations are respectively shown. Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed supporting baseline precisely captured the trend of the lower bound so that all the modes can be filtered out.
Figure S5.8: The gap-elimination course for the found spectrum supporting line with series (45). (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom.
Figure S5.9: The decomposition result of (45) from the subsequent VMD, using the found number of modes and center frequency as initialization. The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are four) and the residual.
Figure S5.10: The iterations for finding the supporting baseline of FFT with respect to time series $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$ are shown. The 1000th(a), 2000th(b), 3000th(c) iterations are respectively shown.
Figure S5.11: The noise-elimination course for the found spectrum supporting line with series $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$. (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
Figure S5.12: (a) and (b)
Figure S5.13: The decomposition result (a) of $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$ from the subsequent VMD, using the found number of modes and center frequency as initialization, compared with original VMD results with IMFS number 2 (b) and 3 (c) respectively. The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are three) and the residual. As compared with (b) and (c), for original VMD, both the results has reconstruction error twice as large as ours.
Figure S5.14: The iteration for finding the baseline of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The 500th(a), 1000th(b), 1500th(c) iterations are respectively shown.
Figure S5.15: The gap-elimination course for the found spectrum supporting line with series $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$. (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
Figure S5.16: The decomposition result of $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$ from the subsequent VMD, using the found number of modes and center frequency as initialization. The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are three) and the residual.

### v2

@@
 Abstract
-Variational Mode Decomposition (VMD) plays an important role in many scientific areas, especially for the area of signal processing.
-Unlike the traditional Fourier paradigm, it makes decomposition of a signal possible without any predefined function basis, which gives unprecedented flexibilities while handling narrow-banded signals of varieties.
-However, determining the number and central frequencies of intrinsic mode functions are still open questions in that few studies has been proposed to give a complete method that can theoretically guarantee the global convergence during decomposition.
-In this article, we propose a globally convergent numerical optimization method based on variational convex optimization to automatically determine the number and central frequency of IMFs without any prior knowledge for narrow banded signals, as long as the spectra of each sub-band are identifiable.
-Our method focuses on finding the support baseline of the spectral function, and further separating the significant frequency band regions above the support baseline.
-Unlike pioneer works that focus on optimizations on complex field, our method achieves obtaining the number of decomposed IMFs and center frequencies in real field by combining variational calculus, convex optimization, and numerical solutions of differential equations in real field and theoretical analysis shows our algorithm is guaranteed terminate to one of the optimum as close as possible.
-Experiments also shows that our algorithm converges quickly and can be used in practical engineering to determine the prior information of the number of intrinsic modes and evaluate initial center frequency as prior information to continue the VMD procedure.
+Automatically determining the number of intrinsic mode functions (IMFs) and their center frequencies in Variational Mode Decomposition (VMD) still remains an open mathematical challenge, which often relies on heuristic prior settings, trial-and-error strategies or complex-field based methods that lack of theoretical guarantees on the convergence.
+In this article, we proposed a novel variational method based on real field, focusing on evaluating the supporting baseline of a signal’s spectrum, to further retrieve the intrinsic mode functions.
+Our method treats automatic extraction of modes as a constrained optimization problem that adversarially maximizes the baseline integral while penalizing its curvature, and transform the problem into iteratively solving a fourth-order boundary value problem via Lagrangian duality.
+Furthermore, we establish the rigorous mathematical proof on the global convergence to our algorithm based on dual ascent in function space.
+Comprehensive numerical experiments on artificial and real-world signals including electrocardiogram (ECG) data show that our method can provide accurate estimates of IMFs and center frequencies, and comparison with methods like Successive VMD also shows our advantages in avoiding redundant modes while ensuring the recovery of necessary components, indicating that we have provided a robust, theoretically grounded initialization routine for VMD.
 keywords:
-narrow-banded signals,variational optimization, Intrinsic Mode Function(IMF)
+narrow-banded signals;variational optimization;Intrinsic Mode Function
 {MSCcodes}
@@
 1 Introduction
-Variational Mode Decomposition (VMD) [5] decomposes original signals into multiple Intrinsic Mode Functions (IMF) with limited bandwidth with the help of variational optimization.
-The core of VMD is to minimize the sum of the bandwidths with respect to the estimated IMFs as well as to ensure the sum of all the IMFs is equal to the original signal.
-The paradigm of VMD is a significant progress in the field of signal processing in that it not only inherits the core of EMD that breaks the traditional Fourier-transform alike framework for signal decomposition where a signal must be decomposed in terms of fixed signal basis like sine and cosine functions, but is also considered as the first trial to solve the signal decomposition problem with the help of variational optimization.
-VMD plays an important role in modern signal decomposition after proposition.
-In the field of mechanics, [14] uses VMD to decompose the vibration signals of planetary gearboxes under variable-speed conditions, in biomedical field, VMD can be used to diagnose the seizure of epileptic by retrieving the feature of decomposed Electroencephalography (EEG) signals [21] and can also help to denoise the Electrocardiogram (ECG) signals [9].
-In the field of Earth Physics signals, it not only helps to increase the Signal-to-Noise Ratio(SNR) by combining with adaptive non-local mean filtering [7], but also helps the prediction of monthly runoff by decomposing the signal into trend, periodic and random components as the input feature of downside neural networks [16].
-Besides, VMD also provides the potential for the channel identification in communications [1] and helps to the visualization of the protein structure [15].
-We show the detail of VMD in Supplementary Materials Section 1.
-Although VMD has provided flexibility in decomposing signals, there are factors that hinder its further applications in that the number of intrinsic mode function must be determined manually as a piror [19, 10, 18, 17] in coordination with penalty factor without a common principle.
-Besides, VMD procedure works in complex field and to our knowledge there is no rigorous analysis on the convergence[5].
-In order to alleviate the problem, the subsequent research either needs the knowledge to the range of the number of modes [11], or retrieves the IMFs recursively [12], or highly dependent on a predefined window-size to retrieve the peak and valley of spectra [6].
-To our best knowledge, although there are some research focusing on wide-band extension of VMD [22], yet even to narrow-band signals, there are almost no complete research that can simultaneously be rigorous in theoretical analysis, while obtaining estimation of the number of modes and center frequency without any prior knowledge to the signal, and achieving non-recursive iterations.
-In this article, we propose a variation-optimization based, globally convergent method to automatically determine the IMFs numbers and the corresponding center frequencies for narrow-banded signals.
-In short, we model the modes retrieving problem as adaptively extracting the supporting baseline of the signal spectrum amplitude function.
-By subtracting the extracted baseline from the signal spectrum amplitude function, we retrieve the significant frequency band components of the signal and extract the various modes of the narrow-band signal.
-The extracted supporting baseline is acquired by adversarially maximizing the integral of the baseline while minimizing the integral of its curvature.
-The supporting baseline is computed by iteratively solving a variational problem that is degenerated into the solution of a normal ordinary differential equation, and under proper condition and weight parameters the method always converges to the optimal curve that can be used to determine the center frequencies and the number of IMFs after proper estimation of the gap residual at the bottom.
-Once the number of IMFs and initialized frequencies are determined by our method, these parameters can be used to start the following VMD procedure.
-Experiments show that our method can quickly get accurate estimation about the number of IMFs and the corresponding initial center frequencies as prior to the following VMD procedure.
-Our contribution can be summarized as 3 folds.
-First, as far as we know, in determining the number and central frequencies of IMFs, we are the first one to not only give a feasible algorithm that can decompose a narrow-banded signal automatically but also prove the global convergence in a rigorous mathematical view.
-Second, our proof generalizes the classical dual ascent method into functionals, which expands the scope of the application for the classical algorithm.
-Third, experiments show that our algorithm can filter out all the potential modes without any prior setting or assumptions to the number of modes, and also show good performance in both precision and speed even when the center frequencies of different modes are very close.
+Variational Mode Decomposition (VMD) decomposes original signals into multiple Intrinsic Mode Functions (IMF) with limited bandwidth with the help of variational optimization.[5] The core of VMD is to minimize the sum of the bandwidths with respect to the estimated IMFs as well as to ensure the sum of all the IMFs is equal to the original signal.
+In VMD, it is assumed the frequency of each mode is almost compact around a center pulsation and the associated analytic signal is computed by means of Hilbert Transform to get the unilateral frequency spectrum.
+Then each mode’s frequency spectrum was shifted into baseband by multiplying a frequency modulation factor with respect to the estimated center frequency, transforming each band-pass mode into low pass, and the bandwidth can be estimated by $H^{1}$ Gaussian smoothness.
+Afterwards, they decompose the original signal into each modes by minimizing the sum of the bandwidth of each mode with the constraints that the summation of all modes is the original signal.
+In short, their key variation optimization can be listed as following:
+(1) $\min\limits_{u_{k},\omega_{k}}\left\{\sum\limits_{k}\left\|\partial_{t}\left[\left(\delta(t)+\frac{\text{j}}{\pi t}\right)*u_{k}(t)\right]e^{-\text{j}\omega_{k}t}\right\|_{2}^{2}\right\}$
+(2) $\text{s.t.}\quad\sum\limits_{k}u_{k}=f$
+so that each mode can be altered into an associated analytical signal by using Hilbert Transform and shifted into base-band for the following estimation of bandwidth using $H^{1}$ Gaussian smoothness with the shifted demodulated components.
+To solve the problem, one can construct an augmented Lagrangian Function by introducing both quadratic penalty and Lagrangian Multipliers as follows:
+(3) $\mathcal{L}(u_{k},\omega_{k},\lambda)$ $=\alpha\sum_{k}\left\|\partial_{t}\left[\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)\right]e^{-j\omega_{k}t}\right\|_{2}^{2}$
+$\quad\quad\quad\quad\quad\quad\quad+\left\|f-\sum u_{k}\right\|_{2}^{2}+\langle\lambda,f-\sum u_{k}\rangle$
+This problem can be solved by using ADMM(Alternate Direction Method of Multipliers), in which each component can be updated by solving the equivalent minimization problem:
+(4) $u_{k}^{n+1}=\underset{u_{k}\in\mathbb{R}}{\arg\min}\left\{\alpha\left\|\partial_{t}\left[\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)\right]e^{-j\omega_{k}t}\right\|_{2}^{2}+\left\|f-\sum u_{i}+\frac{\lambda}{2}\right\|_{2}^{2}\right\}.$
+One can rewrite it in frequency domain form, so that convolution becomes multiplication for the convenience of the calculation in subsequent variational operations:
+(5) $\hat{u}_{k}^{n+1}=\underset{\hat{u}_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\alpha\left\|j\omega\left[1+\text{sgn}(\omega+\omega_{k})\right]\hat{u}_{k}(\omega+\omega_{k})\right\|_{2}^{2}+\left\|\hat{f}-\sum\hat{u}_{i}+\frac{\lambda}{2}\right\|_{2}^{2}\right\}.$
+and by substitution of $\omega\rightarrow\omega+\omega_{k}$ to perform a translation and considering the Hermitian symmetry of spectrum with respect to the real signals, both terms can be transformed into half-space integrals with non-negative frequencies:
+(6) $\hat{u}_{k}^{n+1}=\underset{u_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\int_{0}^{\infty}4\alpha(\omega-\omega_{k})^{2}|\hat{u}_{k}(\omega)|^{2}d\omega+2\int\left(\hat{f}-\sum\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)^{2}d\omega\right\}$
+which can be evaluated by vanishing the first ordered variation of positive frequencies for both signal components and center frequencies
+(7) $\hat{u}_{k}^{n+1}=\left(\hat{f}-\sum_{i\neq k}\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)\frac{1}{1+2\alpha(\omega-\omega_{k})^{2}}$
+(8) $\omega_{k}^{n+1}=\frac{\displaystyle\int_{0}^{\infty}\omega\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}{\displaystyle\int_{0}^{\infty}\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}$
+Since its establishment, VMD is always playing an important role in modern signal decomposition in many fields[14, 21, 9, 7, 16, 15, 1].
+Although VMD has provided flexibility in decomposing signals, there are factors that hinder its further applications in that the number of intrinsic mode function must be determined manually as a prior in coordination with penalty factor without a common principle[19, 10, 18, 17].
+In order to alleviate the problem, the subsequent research either needs the knowledge to the range of the number of modes ,[11] retrieves the IMFs recursively ,[12] or highly dependent on a predefined window-size to retrieve the peak and valley of spectra [6].
+All these methods either require prior or rely on schemes devoid of convergence guarantees, making the results contains spurious or missed modes.
+The absence of a well-posed, convergent paradigm for automatically finding the number of IMFs is a crucial factor that hinders the development of VMD.
+To close this gap, we introduce a fundamentally different approach for finding the IMFs.
+Instead of retrieving modes in an trial-and-error way or recursively extracting modes, we focus on the spectrum amplitude and reformulate the problem by introducing supporting baseline, a lower-enveloping curve of a signal’s spectrum which can support the spectrum as tightly as possible while keeping itself as smooth as possible, and transform the problem for finding IMFs into first evaluating the supporting baseline by adversarially maximizing its integral while penalizing its curvature, then retrieving the separated peaks of the original spectrum above the supporting baseline.
+Our main contribution can be summarized as 4 folds.
+First, we give a new perspective for determining the number of IMFs by introducing supporting baseline of spectrum and solve the problem in real field.
+Second, we derive its optimality conditions, establishing a rigorous equivalence between this variational problem and a fourth-order boundary value problem, and establish rigorous mathematical proof for the global convergence to the course of dual ascent iteration in function space.
+Thirdly, we develop an efficient iteration scheme for the algorithm by converting the corresponding difference equation and boundary conditions into matrix format and expanding compatible rule to the component-wise multiplication between matrices and vectors for Hadamard product.
+Finally, we give experiments on both artificial and real signals, and comparison with SVMD to show the robustness of our algorithm and experiments show that our method can quickly get accurate estimation about the number of IMFs and the corresponding initial center frequencies as prior to the following VMD procedure.
 This article will be organized as follows: In the following section we propose the main framework of our algorithm and give the proof of the convergence.
@@
 In the final section we show our conclusion about the main contribution of the paper.
-2 Methodology
+2 Methodology and Convergence Proof
 In this section we introduce our method for estimating the number of IMF modes as well as center frequencies of each band.
-We first introduce the background of the problem and give a brief introduction about Variational Mode Decomposition, and then we propose the concept of Supporting Baseline of a Function, finally we describe our method for finding the baseline and the whole workflow.
+We first introduce the background of the problem, and then we propose the concept of Supporting Baseline of a Function, finally we describe our method for finding the baseline and give the convergence proof for our method.
+Figure 1: Illustration to the concept of supporting baseline: In (a)-(c) the solid line is $y=10\cos\pi x-10\cos 5\pi x+x^{2}$.
+The dashed line in (a) is a proper supporting baseline since it actually fits the lower bound of the solid line as well as ignoring high frequent details of the solid line.
+The dashed line in (b) has a gap to the bottom of the solid line so it is not suitable to be a good supporting baseline.
+In contrast, the dashed line in (c) fits too tightly to the lower bound of the solid line, capturing too much detail of the solid line so that it can be reckoned as the lower envelope rather than a proper supporting baseline.
 2.1 Supporting Baseline of a Function
 Without loss of generality we first introduce the Supporting Baseline of a function.
-For any function $s(x)>0$, the supporting baseline is a curve that tightly fits the lower bound of the function and reflects the overall trend of the lower bound.
+For any function $f(x)>0$, the supporting baseline is a curve that tightly fits the lower bound of the function while reflects the overall trend of the lower bound.
 Fig.
@@
 1, intuitively the supporting baseline should be tightly close to the lower bound of the function, approaching the infimum of functions as close as possible as well as maintain its own smoothness in order to capture the global trend of the lower bound rather than too much details, as capturing too much details will in turn reduce the robustness for separating the modes since there might be some noise added to the signal in either time or frequency domain.
-Figure 1: Illustration to the concept of supporting baseline: In (a)-(c) the solid line is $y=10\cos\pi x-10\cos 5\pi x+x^{2}$.
-The dashed line in (a) is a proper supporting baseline since it actually fits the lower bound of the solid line as well as ignoring high frequent details of the solid line.
-The dashed line in (b) has a gap to the bottom of the solid line so it is not suitable to be a good supporting baseline.
-In contrast, the dashed line in (c) fits too tightly to the lower bound of the solid line, capturing too much detail of the solid line so that it can be reckoned as the lower envelope rather than a proper supporting baseline.
 2.2 Variational Methods for Finding the Supporting Baseline
 To solve the problem for finding the supporting baseline, we use optimization method mainly based on variation and convex optimization[8, 4].
-Let $f(x)\geq 0$ be a known function, the problem of finding the supporting baseline curve of $f(x)$ is equivalent to finding a function $g(x)$ that tightly approaching the infimum of $f(x)$ while maintaining the global trend of its lower bound.
+Let $f(x)\geq 0$ be a known function, the problem of finding the supporting baseline curve of $f(x)$ is equivalent to finding a function $g(x)$ that tightly approaching the point-wise infimum of $f(x)$ while maintaining the global trend of its lower bound.
 This can be viewed as the following optimization problem:
@@
 (14) $F=\alpha(x)g^{\prime\prime 2}(x)-\beta(x)g(x)+\lambda(x)[g(x)-f(x)]-\mu(x)g(x)$
-then the above Lagrange Slack Function is like the form $J[g(x)]=\int_{\Omega}F(x,g,g^{\prime\prime})\text{d}x$ so that if we assume the variation is zero at both ends of the curve, one can easily know the optimal solution of the above function happen to satisfy the simplified form of Euler-Poisson equation [8] if we let the variation vanish, so it can be written as
+then the above Lagrange Slack Function is like the form $J[g(x)]=\int_{\Omega}F(x,g,g^{\prime\prime})\text{d}x$ so that if we assume the variation is zero at both ends of the curve, one can easily know the optimal solution of the above function happen to satisfy the simplified form of Euler-Poisson equation if we let the variation vanish,[8] so it can be written as
 (15) $F_{g}+\cfrac{\text{d}^{2}F_{g^{\prime\prime}}}{\text{d}x^{2}}=0$
@@
 where 0 stands for the left end and $n$ the right end of the curve.
-To solve the boundary value problem of ODE, we use Finite Difference Method [20].
-Note that for a differential function $g(x)$, it’s derivative of first to fourth order can be expressed approximately as the linear combination of $g(x)$ itself and we can take similar operation to the function $\alpha(x)$ (See Supplementary Materials 2).
-Since now different order of derivative with respect to $\alpha(x)$ and $g(x)$ can be expressed as the linear combination of themselves, one can easily transform (11) into the matrix-form as follows:
+To solve the boundary value problem of ODE, we use Finite Difference Method.[20] Note that for a differential function $g(x)$, it’s derivative of first to fourth order can be expressed as the linear combination of $g(x)$ itself in the corresponding differential format and we can take similar operation to the function $\alpha(x)$ (See Supplementary Materials A), so that different orders of derivative with respect to $\alpha(x)$ and $g(x)$ can be expressed as the linear combination of themselves, one can easily transform (19) into the matrix-form as follows:
 (21) $(2\boldsymbol{A^{(2)}\alpha})\odot(\boldsymbol{G^{(2)}g})+(4\boldsymbol{A^{(1)}\alpha})\odot(\boldsymbol{G^{(3)}g})+(2\boldsymbol{\alpha})\odot(\boldsymbol{G^{(4)}g})=\boldsymbol{\beta-\lambda+\mu}$
-where $\boldsymbol{A}^{(n)}$ and $\boldsymbol{G}^{(n)}$ is the conversion matrices (See Supplementary Materials 2) that transform the nth derivatives of $\alpha(x)$ and $g(x)$ into linear expression of the corresponding function, together with the boundary condition to form full-rank matrices, and $\boldsymbol{\alpha}$, $\boldsymbol{g}$, $\boldsymbol{\beta}$, $\boldsymbol{\lambda}$, $\boldsymbol{\mu}$ is obtained by respectively discretizing the corresponding functions, $\odot$ stands for component-wise multiplication between matrices(See Supplementary Materials 3 for details).
-The conversion matrices $\boldsymbol{A}^{(n)}$ and $\boldsymbol{G}^{(n)}$ share the same structure, in which the last four lines depicts the boundary condition and the other the ODE equation, and $\boldsymbol{g}$, $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$, $\boldsymbol{\lambda}$, $\boldsymbol{\mu}$ are column vectors in which each element is a uniformly discrete sampled value on the function $g(x)$, $\alpha(x)$, $\beta(x)$, $\lambda(x)$, $\mu(x)$, respectively.
-2.3 Proof of the Convergence
-Algorithm 1 The Projection Gradient Descent Method for Evaluating Support Baseline
+[…]
+Algorithm 1 The Key Steps for Evaluating Support Baseline
 1: Input $\{f(x)\}$
@@
 12:until convergence: $\|g^{n}-g^{n-1}\|_{2}^{2}/\|g^{n}\|_{2}^{2}<\epsilon$
-In this section we show the global convergence of the Algorithm 1.
-In our algorithm we have in fact solved a functional Extreme Value Problem with respect to the spectrum to the original signal.
-Despite the fact that the Fourier transform spectrum of a general signal may not necessarily be smooth, yet on one side, as in practical digital signal processing, the signal are discretized and can be smoothly interpolated, indicating the function processed by our algorithm are practically smooth which enable the reasonable use of commonly used assumptions that are advantageous for proving algorithm convergence, and on the other hand, subsequent numerical experiments have also shown that such assumptions not only do not affect the range of signals that can be processed, but are also very helpful for proving convergence.
-Therefore, in our proof, we assume that the spectrum is smooth so that we can extend the Lagrangian dual gradient ascent to functionals.
-Our proof combines functional, variational optimization and ODE theories in the framework of gradient projection, and can be divided into 3 steps.
-First we prove that we are dealing with a convex functional optimization problem, where the feasible domain is also a convex set.
-Next we show that under the condition of our algorithm, the Lagrangian dual gap is zero [4, 2] so that it is possible to find the optimum by dual ascent method with some modification from vectors to functional.
-After that, the convergence of iteration in dual space is shown and it is also shown that the optimal function can also be approached during the iterations by utilizing the dependence of differential equation solutions on coefficients.
-For the sake of limited spaces, we move the proof of convexity and duality into Supplementary Materials 4 and only give the proof to the convergence of iteration in Algorithm 1 in this part.
-We first restate the iteration in Algorithm 1. Notice that during the iteration, we have
+2.3 Proof of the Convergence
+One of the core contribution of this work is to provide a deterministic, global convergence guarantees for the problem of automatic mode detection, which is typically unaddressed in the pioneer works.
+we prove that the algorithm generates sequences converging to an optimal solution of the original problem.
+The proof is built upon several folds:
+Convexity and Duality: Proof of the problem’s convexity and strong duality.
+Well-posedness of Iterations: Demonstration that each iterative step solves a well-posed subproblem with a unique solution.
+Convergence in Dual Space: Impose functional functional to dual ascent method and analyse its convergence under both unique and non-unique optimal dual solutions.
+Convergence in Primal Space: Establish the convergence of the primal sequence via the continuous dependence of solutions on parameters.
+For the conciseness of the proof, we leave the detail of the the convexity and duality into Appendix and focus on the well-posedness of iterations and convergence in dual and primal spaces here.
+2.3.1 Convexity and Duality
+In this part we state the main results about the convexity and duality, and we put the details in Appendix B.1 and B.2.
+Lemma 2.1.
+The objection functional of (10) is convex, and the feasible set is also convex.
+Lemma 2.2.
+The dual of objection functional of (10) is concave, with the dual-gap zero.
+2.3.2 Well-posedness of Iterations
+Before analyzing the convergence of the iterative algorithm, we must ensure that the subproblem solved at each iteration is well-posed.
+Specifically, as in each iteration, our primal solution $g_{k+1}(x)$ is given by dual variables $(\lambda_{k}(x),\mu_{k}(x))$ at the k-th step and we transform the variational problem into a linear fourth-order boundary value problem (BVP), the well-posedness of this BVP encompassing the existence and uniqueness of the solution is crucial to guarantee that our iterative scheme is not only well-defined but also stable against small numerical perturbations.
+Since $\alpha(x)$ and $\beta(x)$ in our algorithm can be considered as weights function that are always positive, we can establish the existence and uniqueness of our solution to (19) below.
+Theorem 2.3.
+Assume $\alpha(x)>0$, then the solution of (19) with $g(0)=0,\\ g(n)=0,g^{\prime}(0)=0,g^{\prime}(n)=0$ is $g(x)\equiv 0$.
+Proof 2.4.
+Recall that when dealing with the homogeneous condition, we exactly have
+(22) $g(x)=\int_{a}^{x}\left(\int_{a}^{x}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t\right)\text{d}t+C_{3}x+C_{4}$
+where
+(23) $g(a)=g(b)=g^{\prime}(a)=g^{\prime}(b)=0$
+It is obvious that $C_{3}$ and $C_{4}$ are zero by $g(a)$ and $g^{\prime}(a)$, and $g^{\prime}(b)=0$ is equivalent to
+(24) $\int_{a}^{b}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t=0$
+Since $\alpha(t)>0$ and $\alpha(t)\in C[a,b]$, since $C_{1}t+C_{2}$ is a linear function so that if the zero point is not in $[a,b]$ then it is not possible to make (24) hold.
+If there exists $t\in[a,b]$ that makes $C_{1}t+C_{2}$ zero, then the upper limit integral function $\Phi(x)=\displaystyle\int_{a}^{x}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t$ has no zero points in $(a,b)$, thus makes $g(b)=\displaystyle\int_{a}^{b}\Phi(t)\text{d}t$ non-zero, which is paradoxical to our assumption if not both $C_{1}$ and $C_{2}$ are zero.
+Then according to Fredholm Alternative Theorem, we can assert that during each iteration, there always exist a unique solution to g(x)g(x)[26, 27, 28].
+Corollary 2.5.
+For each iteration of (21), there always exists a unique solution.
+2.3.3 Convergence in Dual Space
+To facilitate the convergence analysis in the dual space, we introduce the following regularity assumptions on the primal problem:
+The target spectrum function $f(x)$ is Lipschitz continuous on $\mathcal{D}$.
+The solution $g(x)$ to the subproblem and the iterates $g_{k}(x)$ reside in $C^{2}(\mathcal{D})$ with uniformly bounded second derivatives.
+These assumptions are standard for analyzing variational problems involving curvature penalties.
+In practice, for discrete and potentially noisy spectral data, $f(x)$ can be considered as a smooth interpolation of the observed points.
+Crucially, the core object of interest—the supporting baseline—is intrinsically a smooth function due to the curvature penalty term in the objective.
+Therefore, the analysis under these smoothness assumptions captures the essential behavior of the algorithm.
+We begin the proof by first restating the iteration in Algorithm 1. Notice that during the iteration, we have
 (25) $\left\{\begin{aligned} g_{k+1}(x)&=\underset{g(x)}{\arg\min}\mathcal{L}(g(x),\lambda_{k}(x),\mu_{k}(x))\\ \lambda_{k+1}(x)&=\max[0,\lambda_{k}(x)+\theta(g_{k+1}(x)-f(x))]\\ \mu_{k+1}(x)&=\max[0,\mu_{k}(x)+\gamma(-g_{k+1}(x))]\end{aligned}\right$.
@@
 (26) $\int_{\mathcal{D}}\bigl[m(x)-\Pi m(x)\bigr]\bigl[M(x)-\Pi m(x)\bigr]dx\leq 0$
-Proof 2.2.
+Proof 2.7.
 By expanding the equation (26) we should have the following:
@@
 Moreover, if $m(x)$ is in feasible set so that $m(x)=\Pi m(x)$, then (27) will become zero.
-By using Lemma 15, we can prove that in iteration (14) the multiplier $\lambda_{k}(x)$, $\mu_{k}(x)$ always terminate at one of the optimum. To say in detail, we should have the following theorem:
+Now we prove that in iteration (25) the multiplier $\lambda_{k}(x)$, $\mu_{k}(x)$ always terminate at one of the optimum. To say in detail, we should have the following theorem:
 Theorem 2.8.
 Let $\mathcal{G}(\lambda(x),\mu(x))=\underset{g(x)}{\arg\min}\mathcal{L}(g(x),\lambda(x),\mu(x))$ and $\lambda_{k}(x)$, $\mu_{k}(x)$ be equicontinuous for all $k=1,2,3,\cdots$, then if $\mathcal{G}$ satisfies
+(28) $||D[\mathcal{G}(\lambda_{1}(x),\mu_{1}(x))$ $-\mathcal{G}(\lambda_{2}(x),\mu_{2}(x))]||$
 […]
 for any $\lambda_{1}(x)$, $\lambda_{2}(x)$, $\mu_{1}(x)$, $\mu_{2}(x)$, where $D$ stands for the variation, and the step size $\eta<\frac{1}{L}$, then as $k\rightarrow+\infty$, the iteration series of $(\lambda_{k}(x),\mu_{k}(x))$ must terminate at the point as close as possible to one of the optimum $(\lambda^{*}(x),\mu^{*}(x))$ for $\mathcal{G}$.
-Proof 2.4.
+Proof 2.9.
 For conciseness, simplifying the integral by introducing matrix-form expression, where
@@
 $(\lambda_{k+1}(x),\mu_{k+1}(x))$.
-If $\mathcal{G}$ is second-order differentiable and satisfies (17) so that $D^{2}\mathcal{G}$ is bounded with [L,L][-L,L].
+If $\mathcal{G}$ is second-order differentiable and satisfies (2.8) so that $D^{2}\mathcal{G}||$ is bounded with [L,L][-L,L].
 Due to the concavity of $\mathcal{G}$ and the left side of the bound, and combine (41) we get
 […]
-if $\eta=\max\{\theta,\gamma\}=\cfrac{1}{L}<\cfrac{2}{L}$, then $\mathcal{G}$ is a monotonously increasing functional series with upper bound due to its concavity, so that $\mathcal{G}$ converges and the difference between adjacent term of both $\mathcal{G}$ and $(\lambda_{k}(x),\mu_{k}(x))$ is approaching zero.
+if $\eta=\max\{\theta,\gamma\}=\frac{1}{L}<\frac{2}{L}$, then $\mathcal{G}$ is a monotonously increasing functional series with upper bound due to its concavity, so that $\mathcal{G}$ converges and the difference between adjacent term of both $\mathcal{G}$ and $(\lambda_{k}(x),\mu_{k}(x))$ is approaching zero.
 Now we prove $\{g_{k}(x)\}$ is guaranteed to approach one of the optimum when the algorithm terminates even if $\mathcal{G}$ is not strongly concave.
-Due to the concavity the iteration on $\mathcal{G}$ should satisfy [2]
+By introducing $l(\boldsymbol{y};\boldsymbol{x})=\mathcal{G}(\boldsymbol{x})+D\mathcal{G}(\boldsymbol{x})(\boldsymbol{y}-\boldsymbol{x})$ and due to the concavity the iteration on $\mathcal{G}$ should satisfy [2]
 $l\left(\begin{bmatrix}\lambda_{k+1}(x)\\ \mu_{k+1}(x)\end{bmatrix};\begin{bmatrix}\lambda_{k}(x)\\ \mu_{k}(x)\end{bmatrix}\right)-\cfrac{1}{2\eta}\left\|\begin{bmatrix}\lambda_{k+1}(x)-\lambda_{k}(x)\\ \mu_{k+1}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
 $\geq l\left(\begin{bmatrix}\lambda^{*}(x)\\ \mu^{*}(x)\end{bmatrix};\begin{bmatrix}\lambda_{k}(x)\\ \mu_{k}(x)\end{bmatrix}\right)-\cfrac{1}{2\eta}\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
-(35) $+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}$
-since $\mathcal{G}$ is convex. Note that the spectrum of $D^{2}\mathcal{G}$ is larger than L-L and $\eta=\cfrac{1}{L}$ to replace the left side of (35), and use the concavity of $\mathcal{G}$, we have
-$\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}\leq 2\eta(\mathcal{G}(\lambda_{k+1}(x),\mu_{k+1}(x))$
-(36) $-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x)))+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
+(46) $+\cfrac{1}{2\eta}\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}$
+since $\mathcal{G}$ is convex. Due to the concavity of $\mathcal{G}$, we have
+(47) $\mathcal{G}(\lambda_{k+1}(x),\mu_{k+1}(x))\geq l\left(\begin{bmatrix}\lambda_{k+1}(x)\\ \mu_{k+1}(x)\end{bmatrix};\begin{bmatrix}\lambda_{k}(x)\\ \mu_{k}(x)\end{bmatrix}\right)-\cfrac{1}{2\eta}\left\|\begin{bmatrix}\lambda_{k+1}(x)-\lambda_{k}(x)\\ \mu_{k+1}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
+and
+(48) $l\left(\begin{bmatrix}\lambda^{*}(x)\\ \mu^{*}(x)\end{bmatrix};\begin{bmatrix}\lambda_{k}(x)\\ \mu_{k}(x)\end{bmatrix}\right)\geq\mathcal{G}(\lambda^{*}(x),\mu^{*}(x))$
+Substituting (47) and (48) into (46), then we have
+$\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}$ $\leq 2\eta[\mathcal{G}(\lambda_{k+1}(x),\mu_{k+1}(x))$
+(49) $-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x))]+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
 Sum up (49) for each $k$ and eliminate the repeated terms that occur at the both ends of the inequality, we have
-$\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}\leq 2\eta\sum\limits_{i=0}^{k}(\mathcal{G}(\lambda_{i+1}(x),\mu_{i+1}(x))$
-(37) $-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x)))+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{0}(x)\\ \mu^{*}(x)-\mu_{0}(x)\end{bmatrix}\right\|_{2}^{2}$
+$\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}$ $\leq 2\eta\sum\limits_{i=0}^{k}[\mathcal{G}(\lambda_{i+1}(x),\mu_{i+1}(x))$
+(50) $-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x))]+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{0}(x)\\ \mu^{*}(x)-\mu_{0}(x)\end{bmatrix}\right\|_{2}^{2}$
 As the left side is non-negative, and recall in (45) we know that $\{\mathcal{G}_{k}\}$ is non-decreasing and $\mathcal{G}^{*}$ is the supreme of $\mathcal{G}$, hence $\mathcal{G}(\lambda_{k+1}(x),\mu_{k+1}(x))-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x))$ is the largest among all the terms involved in the summation, by applying reasonable inequality scaling we must have
@@
 which shows the iteration is approaching the optimum. From (49) and by utilizing the monotonicity of the functional $\mathcal{G}$ during iteration we have
-(39) $\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}<\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
+(52) $\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}\leq\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
 which shows the distance between the iterating points and the optimum is always decreasing, since the distance is always non-negative so that it has lower bound, $(\lambda_{k}(x),\mu_{k}(x))-(\lambda^{*}(x),\mu^{*}(x))$ is a convergent sequence so that $(\lambda_{k}(x),\mu_{k}(x))$ is bounded.
-According to Arzelà-Ascoli Theorem, there exists some sub-sequences $(\lambda_{k_{i}}(x),\mu_{k_{i}}(x))$ that converges, so that we cannot assert that the sequence is necessarily convergent in all cases and need careful discussion.
-One one hand, if there is only one optimum in $\mathcal{G}$, then it is trivial to show that all these sub-sequences converges to the unique optimum by the constraints in (38).
-If $\mathcal{G}$ has more than one optimum, then suppose there exists more than one sub-sequence, for example, $(\lambda_{k_{i}}(x),\mu_{k_{i}}(x))$, $(\lambda_{k_{j}}(x),\mu_{k_{j}}(x))$ that converges to $(\lambda^{*}_{i}(x),\mu^{*}_{i}(x))$, $(\lambda^{*}_{j}(x),\mu^{*}_{j}(x))$ respectively, then due to (34) we show that although the sequence might have converged subsequence that approaches to different optimum, the adjacent term tends to zero so that the adjacent primal solution in iteration also tends to zero.
-On the other hand, as numerical algorithms will introduce error inevitably, if after some iteration $(\lambda_{k}(x),\mu_{k}(x))$ is exactly the unique optimum so that the right hand of (39) is zero, then from monotonicity of iteration we know $(\lambda_{k+1}(x),\mu_{k+1}(x))=(\lambda_{k}(x),\mu_{k}(x))$.
-Even if $\mathcal{G}$ has more than one optimum and after some iteration $(\lambda_{k}(x),\mu_{k}(x))$ approaches one of the optima $(\lambda^{*}_{1}(x),\mu^{*}_{1}(x))$, since there is another optimum $(\lambda^{*}_{2}(x),\mu^{*}_{2}(x))$, then we have
+According to Arzelà-Ascoli Theorem, we can only assert that there exists some sub-sequences $(\lambda_{k_{i}}(x),\mu_{k_{i}}(x))$ that converges, which does not indicate that the whole iterating sequence is necessarily convergent in all cases and needs careful discussion.
+There are four cases when the iterating sequence is approaching the optimum:
+CASE I: If there is only one optima in $\mathcal{G}$, then it is trivial to show that all these sub-sequences converges to the unique optimum by the constraints in (51).
+CASE II: If the optimum of $\mathcal{G}$ forms a convex set, and during the iteration the sequence $(\lambda_{k}(x),\mu_{k}(x))$ is approaching and never fall into the optimal set, then from the middle parts of (51) and (45) we know that the value of $\mathcal{G}$ is approaching the minima and the distance between adjacent term of the sequence $(\lambda_{k}(x),\mu_{k}(x))$ and $(\lambda_{k+1}(x),\mu_{k+1}(x))$ is yielding zero.
+Thus we know in this case the iterating sequence in dual space can always terminate to the point as close as possible to one of the optimum, although the classical convergent behaviour for the sequence in this case might not be guaranteed.
+CASE III: If the iterating sequence has approached to the unique optima during some steps, then due to (52) the next iteration will still stay at the unique optima.
+CASE IV: If the optimum of $\mathcal{G}$ forms a convex set, and the iterating sequence has approached to one of the optimum $(\lambda^{*}_{1}(x),\mu^{*}_{1}(x))$ early during some iterating step, then since in the convex set there are at least another optima $(\lambda^{*}_{2}(x),\mu^{*}_{2}(x))$, then we have
 (53) $\mathcal{G}(m\lambda^{*}_{1}(x)$ $+(1-m)\lambda^{*}_{2}(x),m\mu^{*}_{1}(x)+(1-m)\mu^{*}_{2}(x))$
@@
 thus the variation direction is always orthogonal to any lines connected between different optima, as a consequence, when projected again during the next iteration, $(\lambda_{k+1}(x),\mu_{k+1}(x))$ must overlap with $(\lambda_{k}(x),\mu_{k}(x))$, so we have $(\lambda_{k+1}(x),\mu_{k+1}(x))=(\lambda_{k}(x),\mu_{k}(x))$.
-In all cases we show $(\lambda_{k}(x),\mu_{k}(x))$ is always convergent to at least one of the optimum and stay still as long as it arrives the optimum during the iteration even if there are more than one optimum in $\mathcal{G}$.
-Now we turn to prove the convergence of series $g_{k}(x)$.
-As each $g_{k}(x)$ is obtained by solving an ODE function as the optimum with respect to a variational problem, we should first prove the existence and uniqueness to the solution of $g$ for each iteration.
+In all cases we show $(\lambda_{k}(x),\mu_{k}(x))$ can always terminate at the point as close as possible to at least one of the optimum and stay still as long as it arrives the optimum during the iteration even if there are more than one optimum in $\mathcal{G}$.
+2.3.4 Convergence in Primal Space
 Theorem 2.10.
-The solution of (11) with $g(0)=0,g(n)=0,g^{\prime}(0)=0,g^{\prime}(n)=0$ is $g(x)\equiv 0$.
-Proof 2.6.
-Note that the differential equation of the boundary value problem we are solving has only zero solutions under the corresponding homogeneous differential equation and homogeneous boundary conditions, in order to prove the existence and uniqueness of our solution in each iteration [3].
-Recall that when dealing with the homogeneous condition, we exactly have
-(42) $g(x)=\int_{a}^{x}\left(\int_{a}^{x}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t\right)\text{d}t+C_{3}x+C_{4}$
-where
-(43) $g(a)=g(b)=g^{\prime}(a)=g^{\prime}(b)=0$
-It is obvious that $C_{3}$ and $C_{4}$ is zero by $g(a)$ and $g^{\prime}(a)$, and $g^{\prime}(b)=0$ is equivalent to
-(44) $\int_{a}^{b}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t=0$
-If $\alpha(t)>0$ and $\alpha(t)\in C[a,b]$, since $C_{1}t+C_{2}$ is a linear function so that if the zero point is not in $[a,b]$ then it is not possible to make (44) hold.
-If there exists $t\in[a,b]$ that makes $C_{1}t+C_{2}$ zero, then the upper limit integral function $\Phi(x)=\displaystyle\int_{a}^{x}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t$ has no zero points in $(a,b)$, thus makes $g(b)=\displaystyle\int_{a}^{b}\Phi(t)\text{d}t$ non-zero, which is paradoxical to our assumption if not both $C_{1}$ and $C_{2}$ is zero.
-As a consequence, during each iteration, there always exist a unique solution to $g(x)$, so that we can simply discuss the convergence based on the continuous-dependency between the solution and the coefficients.
-Corollary 2.7.
-For each iteration of (13), there always exists a unique solution.
-Now we begin to prove the convergence of $g_{k}(x)$. For this property during the iteration in Algorithm 1, we have the following theorem:
-Theorem 2.8.
-Let $f(x)$ be first-ordered Lipschitz-continuous function that satisfies $|f^{\prime}(x)-f^{\prime}(y)|\leq L|x-y|$ for any $x,y$, and suppose $g_{0}(x)$ satisfies $0\leq g_{0}(x)\leq f(x)$.
-If $\{g_{k}(x)\}$ satisfies
 […]
-[…]
-Then if the step $\eta$ is sufficiently small so that $\eta<\frac{1}{L}$ for each iteration, $g_{k}(x)$ must terminate at one of the optimum.
-Proof 2.9.
-Based on lemma 15, theorem 2.3 and theorem 2.5 we have shown that our original problem has unique solution during each iteration.
-As (11) is a linear ODE, its solution is continuously dependent on its coefficients, so we have $\lim_{k\rightarrow+\infty}g_{k}(x)=g^{*}(x)$ when $\{(\lambda_{k}(x),\mu_{k}(x))\}$ converges or falls early in one of the optima, and
-[…]
-$\{(\lambda_{k}(x),\mu_{k}(x))\}$ has sub-sequences approaching different optimum respectively.
+Proof 2.11.
+As the solution of (10) is continuously dependent on its coefficients, so we have $\lim_{k\rightarrow+\infty}g_{k}(x)=g^{*}(x)$ when $\{(\lambda_{k}(x),\mu_{k}(x))\}$ when in Case I, III and IV, and $g_{k}(x)$ is approaching $g^{*}(x)$ when $g_{k+1}(x)-g_{k}(x)||$ is approaching zero in Case II respectively.
 In all cases we can terminate algorithm and guarantee the $\{g_{k}(x)\}$ terminates to one of the optimum as close as possible.
 3 Implementation
-3.1 Algorithm
-In this part we state the whole implementation of our method.
-Our implementation can be mainly divided into three parts.
-Given a signal to be decomposed, we first use variational method in terms of iteratively constructing such a function sequence that each term of it is the optimal solution obtained by vanishing the variation of the Lagrangian Slack Function under the current evolutionary state with respect to inequality constraint conditions and then updating the weight of constraint using dual ascent method.
-Thirdly, we apply Kernel Density Estimation (KDE) to the amplitude values of each point in the spectrum to eliminate residual gaps at the bottom, which is based on the observation that the residual gap at the bottom will results in a clustering behaviour around the corresponding value that the residual gap share.
-The proper threshold to eliminate the gap for further evaluating the number of IMFs should locate where the value of the KDE function achieves the maximum.
-In short, our whole workflow can be summarized as in Algorithm 2.
+In this chapter, we focus on the detail about the implementation and numerical scheme about the algorithm for evaluating problem (10).
+Our workflow consists of 4 interconnected stages: (1) We discretize (19) by using Finite Difference Method,[20] transforming it into a linear system.
+(2) Crucially, by expressing the derivations of both solution $g(x)$ and the weight function $\alpha(x)$ as linear combination of their nodal values, we recast the entire ODE into a compact matrix form by introducing component-wise multiplication with a mild extension of Hadamard production.
+By imposing well-posed boundary conditions, we make the coefficient matrices full-rank and the whole problem invertible so that the evaluation can be significantly accelerated by modern software toolkits via matrix operation.
+(3) To further accelerate the convergence and make algorithm effective when handling high or low-passed signals, a smooth extrapolation technique is applied to the input spectrum at the both ends.
+(4) Finally, after obtaining the optimal baseline, a statistically grounded threshold determined via kernel density estimation (KDE), is employed to automatically identify significant spectral peaks, thereby completing the mode extraction.
+The following subsections detail each of these components.
+3.1 Spatial Discretization Scheme
+We discretize the solution domain $\Omega=[a,b]$ into $N+1$ equally spaced nodes $x_{0},x_{1},\dots,x_{n}$, where $x_{0}=a$ and $x_{n}=b$, with the grid spacing $h=\dfrac{b-a}{N}$, thus the value of any continuous function $p(x)$ at the nodes are recorded as a column vector $\boldsymbol{p}=[p(x_{0}),p(x_{1}),\dots,p(x_{N})]^{\top}\in\mathbb{R}^{N+1}$.
+Then we employ Five-Point Central Difference Scheme to transform different orders of derivatives into the linear combination with respect to the primal nodal values.
+For instance, for the fourth-order derivative $g^{(4)}(x)$ at interior nodes $x_{i},i=2,3,\dots,N-2$ we have
+(55) $g^{(4)}(x_{i})\approx\cfrac{g(x_{i+2})-4g(x_{i+1})+6g(x_{i})-4g(x_{i-1})+g(x_{i-2})}{h^{4}}$
+and the corresponding boundary conditions can be approximately expressed as
+(56) $\left\{\begin{aligned} g(0)&=g(x_{0})\\ g(n)&=g(x_{N})\\ g^{\prime}(0)&\approx\cfrac{g(x_{1})-g(x_{0})}{h}\\ g^{\prime}(n)&\approx\cfrac{g(x_{N})-g(x_{N-1})}{h}\end{aligned}\right$.
+Note that (55) and (56) can be transformed into compact matrix form via a pentadiagonal matrix $\boldsymbol{G}^{(4)}\in\mathbb{R}^{(N+1)\times(N+1)}$, thus we have
+(57) $\boldsymbol{G}^{(4)}=\begin{bmatrix}1&-4&6&-4&1&0&\cdots&\cdots&\cdots&0\\ 0&1&-4&6&-4&1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&1&-4&6&-4&1&0\\ 0&0&\cdots&\cdots&0&1&-4&6&-4&1\end{bmatrix}$
+thus $\boldsymbol{G}^{(4)}\boldsymbol{g}$ is the fourth-ordered derivative of $\boldsymbol{g}$ at each node and the non-zero entry stands for the corresponding coefficients of difference and boundary conditions.
+$\boldsymbol{G}^{(3)}$, $\boldsymbol{G}^{(2)}$, $\boldsymbol{G}^{(1)}$ and $\boldsymbol{A}^{(4)}$ to $\boldsymbol{A}^{(1)}$ have the similar structure as $\boldsymbol{G}^{(4)}$, we give the detail of each pentadiagonal matrix in A.
+By integrating both the difference strategy and the boundary condition (19) is transformed into (21), where $\odot$ stands for component-wise multiplication.
+However, in (21) $\boldsymbol{g}$ is inseparable so that we cannot resort to effective evaluation.
+To solve this, our $\odot$ is just an extension of standard Hadamard production by adding some compatible broadcasting rules in dealing with column vector, that is $\forall\boldsymbol{A}_{m\times n},\boldsymbol{v}_{m\times 1}$, we have
+(58) $\boldsymbol{v}\odot\boldsymbol{A}=\boldsymbol{A}\odot\boldsymbol{v}=\boldsymbol{A}*\boldsymbol{V}_{m\times n}$
+where * stands for standard Hadamard production and each column of $\boldsymbol{V}$ is just a copy of $\boldsymbol{v}$.
+By applying this rule we have $(\boldsymbol{A}\boldsymbol{a})\odot(\boldsymbol{B}\boldsymbol{b})=((\boldsymbol{A}\boldsymbol{a})\odot\boldsymbol{B})\boldsymbol{b}$, we can transform (21) into
+(59) $((2\boldsymbol{A^{(2)}\alpha})\odot\boldsymbol{G^{(2)}})\boldsymbol{g}+((4\boldsymbol{A^{(1)}\alpha})\odot\boldsymbol{G^{(3)}})\boldsymbol{g}+((2\boldsymbol{\alpha})\odot\boldsymbol{G^{(4)}})\boldsymbol{g}=\boldsymbol{\beta}-\boldsymbol{\lambda}+\boldsymbol{\mu}$
+so that we can treat (21) as common matrix equation and get the solution by simply using matrix inversion with modern mathematical software tookits.
+Figure 2: Iterations for finding the supporting baseline without ((a) and (b)) /with ((c) and (d)) smooth extrapolation to the spectrum when iterated to 500 ((a) and (c)) and 1500 ((b) and (d)) steps for the same spectrum.
+The solid line is the original spectrum and the dotted line is the evaluated supporting baseline.
+With smooth extrapolation to the original spectrum, the speed of convergence has significant improvement.
+3.2 Treatment of Boundary Conditions
+The primal problem requires the supporting baseline $g(x)$ confirming with the value of $f(x)$ and $f^{\prime}(x)$ at both ends as boundary conditions, this conducts linear constraints directly onto the first and last elements to the solution vector of $\boldsymbol{g}$.
+For low and high pass signals of which the spectrum value are high and drop abruptly at either of the ends, this will make $g(x)$ inherits the trend of $f(x)$ at both ends at some interval so that convergence in this case will be very slow with reasonable grid space.
+To alleviate this effect, we adopt a smooth extrapolation strategy by symmetrically appending gently decaying segments at both ends to make the extended value align with the minimum of $f(x)$, and the extended derivative smoothly approach to zero within the extended region.
+This strategy significantly improves the condition number of the coefficient matrix of the discrete system, ensuring the numerical stability of the iterative solution, without any altering the primal spectrum within the original interval.
+Fig.
+2 shows the illustrations.
+3.3 Postprocessing: From Baseline to Mode Identification
+Upon the above course when the algorithm converges, we obtain the optimal discrete baseline vector $\boldsymbol{g}^{*}$ above which significant intrinsic modes corresponds to the parts of the original spectrum $\boldsymbol{f}$ that protrude distinctly.
+However, there are some residual background fluctuations at the bottom that hinders one to objectively distinguish those modes.
+To solve this problem, we assume that the residual consists of two components: first, a background fluttering component originate from noise or minor undulations, of which the amplitude concentrates within a very small range; second, significant protrusions from original modes, which are larger compared with the former and more scattered.
+In this view, when investigating the value of the residual spectrum value distribution, we can intuitively find that most values concentrate near the level where fluttering component resides.
+This inspires us to employ Kernel Density Estimation (KDE) to non-parametrically model the spectrum value distribution as a probability density function:
+(60) $\phi(t)=\cfrac{1}{mh}\sum\limits_{i=1}^{m}K\left(\cfrac{t-r_{j}}{h}\right)$
+where $K(\cdot)$ is Gaussian function, with $h$ the bandwidth and $r_{j}$ the discrete nodes in the residual vector $\boldsymbol{r}$, and we evaluate the best threshold to eliminate the fluttering noise as
+(61) $t^{*}=\underset{t}{\arg\max}\phi(t)$
+so that we can isolate each mode in the function $f(x)-g^{*}(x)-t^{*}$.
+Figure 3: Illustration of the non-consistence of the error between the converged optimal supporting baseline and the original function.
+Fig.
+3a shows the converged $g(x)$ and Fig.
+3b shows f(x)g(x)f(x)-g(x).
+Note that since $g(x)$ does have gap to the lower bound of $f(x)$ so that the bottom of f(x)g(x)f(x)-g(x) is almost not on a horizontal line everywhere, which indicates there are some non-zero gaps.
+These non-zero gaps will have a significant impact on the stability of the subsequent modal number solution.
+3c shows the estimate kernel density pertaining the gap at the bottom.
+3d shows the estimated f(x)g(x)f(x)-g(x) after removal of the gap at the bottom.
+3.4 The Whole Workflow
+Based on the above discrete framework, each step of Algorithm 1 can be implemented by specific linear algebraic operations.
+The whole workflow can be summarized as Algorithm 2.
+We first input the original spectrum $f(x)$, then we initialize weight functions $\alpha(x),\beta(x)$, penalty functions $\lambda(x),\mu(x)$ and the supporting baseline function $g(x)$.
+After that we smoothly extend $f(x)$ at both ends and begin Algorithm 1 until it converges.
+Then we do Kernel Density Estimation to figure out the threshold to eliminate the fluttering noise, so as to isolate band components.
 Algorithm 2 The Whole Workflow
@@
 2: Initialize $\alpha(x),\beta(x),\lambda(x),\mu(x),g(x)$
-3: Pre-process by extending $f(x)$ at both ends.
+3: Smoothly Extending $f(x)$ at both ends.
 4:repeat
@@
 12: Output $\{u_{k}\}$
-3.2 Pre-processing
-In order to solve the equation effectively, before we start the iteration, we first normalize the signal into $[0,1]$, in both domain and range.
-Since in each iteration, our optimal baseline should satisfy the boundary condition of the ODE, so that when dealing with high-pass or low-pass signals where spectrum has sharp slope at the ends, the to-be-evaluated baseline will also exhibits the sharp slope, which will in turn have bad effect on the convergence at the ends.
-To overcome the boundary effect, we extend our original function by making a gentle decay at both ends of the spectrum that not only completes the peaks at both ends but also alleviate the sharp slope effect.
-We will discuss this in detail later in our experiments.
-3.3 Post-processing
-Despite the fact that the iteration procedure to find the supporting baseline converges in reasonable time in most cases when applied with proper extension to the original spectrum of the signal, the shape of the spectrum do have great impact on the convergence rate.
-For example, since our optimization strategy discourage the curvature (via second-order derivative) of the curve, the optimal baseline is prone to converge more faster at points where original spectrum is flat.
-On the other side, although $g(x)$ will eventually adhere closely to the lower bound of $f(x)$, it is difficult to guarantee that the error between $f(x)$ and $g(x)$ calculated after stopping the iteration is consistent at every point so that f(x)g(x)f(x)-g(x) will have some non-zero points where it should be zero, which makes further retrieving modes of difficulty.
-Since $g(x)$ converges to the bottom of $f(x)$, it is reasonable to assume that the gap is typically small so that f(x)g(x)f(x)-g(x) is almost zero except for those frequency points corresponds with effective signal components.
-In order to eliminate the gap, we use Kernel Density Estimation(KDE) method to estimate the best threshold, so that the gap of f(x)g(x)f(x)-g(x) will be exposed where the value of the KDE function achieves the maximum.
-Fig.
-2 illustrates such a phenomenon.
-Figure 2: Illustration of the non-consistence of the error between the converged optimal supporting baseline and the original function.
-Fig.
-2a shows the converged $g(x)$ and Fig.
-2b shows f(x)g(x)f(x)-g(x).
-Note that since $g(x)$ does have gap to the lower bound of $f(x)$ so that the bottom of f(x)g(x)f(x)-g(x) is almost not on a horizontal line everywhere, which indicates there are some non-zero gaps.
-These non-zero gaps will have a significant impact on the stability of the subsequent modal number solution.
-2c shows the estimate kernel density pertaining the gap at the bottom.
-2d shows the estimated f(x)g(x)f(x)-g(x) after removal of the gap at the bottom.
 4 Numerical Results
-In this section we first show our experiments to our method.
-In order to prove the effectiveness of our method, we choose some standard function that were repeatedly introduced as benchmarks in some pioneer works [5, 13, 12], yet pay more attention on investigating the evolution of our algorithm during the course of iteration, to show the convergence and robustness of our method.
-Since our algorithm can be reckoned as an algorithm for finding the number of intrinsic modes and initial center frequencies that the proceeding VMD module needs, our experimental focus is on exploring the behaviour of the algorithm towards various signals in the process of finding the baseline and talking about the KDE process for eliminating the gap, rather than elaborating the performance brought by the subsequent VMD decomposition in the final decomposition results.
-Then we turn to discuss why we should and how we eliminate the boundary effect by extending the spectrum obtained from the FFT of original signals.
+In this section, we show some experiments to validate the applicability of our methods.
+Our experiments focus on 3 goals: (1) To validate the convergence and robustness of our method under synthetic signals like single-modality signal, dual-modality signal, segmented signal, comb-spectrum signal, dense-modality signal to demonstrate the extensive adaptability of the model across modes with various distribution.
+(2) To compare our method against the most challenge-able method Successive Variational Mode Decomposition (SVMD), another recursive, adaptive method determining the intrinsic modes, exhibiting the advantage of our method on not generating redundant modes or eliminating necessary modes.
+(3) To understand the applicability of our method in real-world signals like electrocardiogram (ECG) signals, showing the accuracy of decomposition by investigation of the signals reconstructed from the decomposed modes.
+All experiments are conducted under the same parameter settings unless otherwise stated.
+The stopping criterion is set as $\dfrac{\|g_{n}-g_{n-1}\|}{\|g_{n}\|}<10^{-4}$.
+The number of discrete grid points is fixed at $N=200$ for the normalized frequency domain $[0,1]$.
+Detailed descriptions of the test signals and comparative protocols are provided in the corresponding subsections.
+Figure 4: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) with respect to the error for finding the optimum supporting baseline of FFT with respect to time series $y(t)=100\sin 20\pi t$.
+The iteration lasts for 5161 steps and costs 1.88 seconds.
 4.1 Experiments on Finding the Supporting Baseline
-In this part we show our experiments on finding the supporting baseline for some benchmark signals that were frequently used to test the performance of algorithms in several articles [5, 12].
-As default, all signals are obtained by linearly normalizing the domain of the original function to between 0 and 1, and the range of which is normalized into 0 and 1 before iterations to keep the stability.
-We uniformly add 200 points to the domain and only preserve half of the transformed spectrum as our initial spectrum to be iterated according to the properties of Fourier Transform with respect to real signals.
-In this way, the Nyquist frequency is 100Hz, which means the highest signal angular frequency that can be effectively represented is $200\pi$, which corresponds to 1 in our normalized FFT results.
-Our convergence criterion is $\cfrac{\|g_{n}-g_{n-1}\|}{\|g_{n}\|}<10^{-4}$.
 We choose five typical signals including single mode signal, multi-mode signal, piecewise continuous signal, signal that are not necessarily narrow-banded, composite signal in which the center frequency of its components are very close.
-Due to the space limitations, we only illustrate the final convergent state of the iteration for the evaluation of the support baseline, as well as the convergent curve with respect to the error, and indicate the number of iteration steps in the caption.
-The intermediate iteration process status and the decomposed signal is given in detail in the supplementary materials 5.
+We only illustrate the initial and final state of the iteration for the evaluation of the support baseline, as well as the convergent curve with respect to the error in logarithm to indicate the number of iteration steps in the caption, and leave the detail of decomposition result in C.
 Experiment 1: The first example is a single mode signal, on the other words, the signal is narrow-banded and only contains one center frequency.
@@
 As we see the normalized center frequency is $0.1$.
-Figure 3: The initial, final convergent state and convergence curve in logarithmic coordinate system with respect to the error for finding the optimum supporting baseline of FFT with respect to time series $y(t)=100\sin 20\pi t$.
-The iteration lasts for 5161 steps and costs 1.88 seconds.
+The algorithm takes 5161 steps and 1.88 seconds to converge.
 Experiment 2: The second example is a multi-mode signal, $y(t)=10\cos(10\pi t)+20\sin(20\pi t)$, which has two modes, exactly.
@@
 However, our algorithm can still survive where center frequencies of different modes are very close, since from 6000th iteration and on, the support baseline forms a small bulge at the close frequency points, which just separates the two frequency points.
-Figure 4: The final convergent state in logarithmic coordinate system for finding the supporting baseline of FFT with respect to FFT of $y(t)=10\cos 10\pi t+20\sin 20\pi t$.
+Figure 5: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the supporting baseline of FFT with respect to FFT of $y(t)=10\cos 10\pi t+20\sin 20\pi t$.
 The iteration lasts for 5258 steps and costs 1.89 seconds.
@@
 To our surprise, the iteration converges even more quickly than those signals that has fewer modes.
-In our cases, it takes about 15000 iterations to achieve the convergence.
-Figure 5: The final convergent state in logarithmic coordinate system for finding the supporting baseline of FFT with respect to time series (45).
+In our cases, it takes about 3026 iterations to achieve the convergence.
+Figure 6: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the supporting baseline of FFT with respect to time series (62).
 Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed supporting baseline precisely captured the trend of the lower bound so that all the modes can be filtered out.
@@
 As we see that, the spectrum of the signal exhibits two main peaks, however the signal is not a completely narrow-band signal since it exhibits comb-shaped spectrum with respect to harmonic components, although the energy of those harmonic components are relatively very small(Fig.
-6).
+7).
 Our algorithm extracts three components of the signal, despite that in the mainstream view there should be only two IMFs.
 However, since it is not a narrow-band signal, the difficulty of decomposition objectively exists since one can not even provide a specific standard to determine whether these small harmonic components should be considered as an intrinsic mode.
-We also observe that, even in [5, 12], when carefully zoomed in, there does exist small oscillations in the first decomposed component which indicates the non-purity of it.
+We also observe that, even in [5, 12] when carefully zoomed in, there does exist small oscillations in the first decomposed component which indicates the non-purity of it.
 Despite the fact that our results has three components, every separate component exhibits better purity than the pioneer works.
-On the other side, from the perspective of reconstruction error, it is shown in this example that, results generated by algorithm in [5] with both 2 and 3 modes has reconstruction error as twice as large as ours (0.1 ours vs 0.2 theirs, see S5.13a), which also indicates that our algorithm has better performance in this case.
-Figure 6: The final convergent state in logarithmic coordinate system for finding the baseline of function $\frac{1}{1.2+\cos 2\pi t}+\frac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$.
+On the other side, from the perspective of reconstruction error, it is shown in this example that, results generated by algorithm in [5] with both 2 and 3 modes has reconstruction error as twice as large as ours, which also indicates that our algorithm has better performance in this case.
+Figure 7: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the baseline of function $\frac{1}{1.2+\cos 2\pi t}+\frac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$.
 The iteration lasts for 3463 steps and costs 1.25 seconds.
+Figure 7: The decomposition result (a) of $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$ from the subsequent VMD, using the found number of modes and center frequency as initialization, compared with original VMD results with IMFS number 2 (b) and 3 (c) respectively.
+The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are three) and the residual.
+As compared with (b) and (c), for original VMD, both the results has reconstruction error twice as large as ours.
 Experiment 5: Next we show even when the central frequencies are very close, our algorithm can also work well.
@@
 The signal has one low-frequency component and 10 alternating current components with different amplitudes, of which the minimum difference between frequency gaps is only 0.04 after normalization and our algorithm can still survive in figuring out all the components.
-Figure 7: The final convergent state in logarithmic coordinate system for finding the baseline of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The iteration lasts for 1768 steps and costs 0.64 seconds.
-4.2 Experiments on Extension of Original Spectrum at Both Ends
-Although supporting baseline can be retrieved by the method stated above, in real cases there are circumstances where convergence is very slow.
-For example, since the baseline must satisfy the boundary condition on both side, the trend of spectrum at both ends has a great impact to the shape of the iterating curve.
-In particular, for low or high pass signals of which the spectrum have a significant slope trend in low and high frequency regions, the iterating curve is prone to adhere the slope at both ends and continue to keep such inertia to some extent even when slipping out of those slope regions.
-This phenomenon brings us two bad effects in analysing signals: On one side, for low or high pass signals, their spectrum at both ends always exhibit half peaks since the scope of representable spectrum from FFT is always limited and the spectrum of those signals will inevitably hits the boundary, making the peak not complete and the estimation of center frequency inaccurate.
-On the other side, inertia effect will be harmful to the shape of the baseline, since in regions adjacent to low and high frequency bands, the function value of supporting baseline will have very large gap to the lower bound of the original spectrum due to the inertia of the downward trend.
-In order to circumvent such effect, we extend the original spectrum by forcing the spectrum to gently drop to the minimum of original spectrum and leave some extra flat intervals at both ends.
-This method not only take the advantage of trend effect, in that the curve will tend to keep horizontally at both ends which can make the peak be finally filtered out, but also complete the peaks at both ends without any harmful effects to the other regions as the experiment shows.
-Fig.
-8 shows the evolution of spectrum baseline during the iteration with respect to the signal $y(t)=60t+60\cos 200\pi t$.
-This typical signal is a band-stop signal which shares both high and low pass feature.
-Note that if we don’t extend the spectrum, the baseline exhibits a sharp decreasing in the low frequency area and can not catch up the lower bound in the interval $[0.05,0.4]$ and the similar phenomenon also happens at the high-frequency area, even after 1500 iterations.
-This shows the necessity of proper extension to the original spectrum.
-Figure 8: Iterations for finding the supporting baseline without extending the spectrum to the signal $y(t)=60t+60\cos 200\pi t$.
-The 500th(a), 1500th(b) iterations with no extension and the 500th(c), 1500th(d) iterations with extension are shown respectively.
-The solid line are the original spectrum and the dotted line are the evaluated supporting baseline during the iterations.
+Figure 8: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the baseline of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The iteration lasts for 1768 steps and costs 0.64 seconds.
 4.2 Experiments on Central Frequencies
@@
 As shown in Table 1, in most cases, our estimation to the center frequencies has low difference with respect to the pure VMD.
-It can be found that, in Experiment 4, the error between estimated center frequencies of ours and pure VMD’s are larger than the average, however, as analysed in Experiment 4, on one side, the original VMD, no matter given IMFs number 2 or 3, fails to achieve reconstruction error as small as ours, on the other side, as shown in Fig.S5.13a(c) even when both started at IMFs number 3, the original VMD seems produce replicated modes at the Gaussian-alike component, with almost the same amplitude and shape, while ours provide relatively independent components when given our initial estimation for the center frequencies, which seems to indicates that our estimation in this case is more reliable.
+It can be found that, in Experiment 4, the error between estimated center frequencies of ours and pure VMD’s are larger than the average, however, as analysed in Experiment 4, on one side, the original VMD, no matter given IMFs number 2 or 3, fails to achieve reconstruction error as small as ours, on the other side, as shown in Fig.7c even when both started at IMFs number 3, the original VMD seems produce replicated modes at the Gaussian-alike component, with almost the same amplitude and shape, while ours provide relatively independent components when given our initial estimation for the center frequencies, which seems to indicates that our estimation in this case is more reliable.
+Figure 9: The original spectrums for signal $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$.
+Figure 10: The decomposed spectrums for signal $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$ by our method.
+4.3 Experiments Compared with Variational Mode Decomposition
+We noticed that, the Successive Variational Mode Decomposition(SVMD) method is another competitive method that can also automatically determine the number of IMFs for VMD[12] without any prior knowledge although it also lacks of rigorous proof on the convergence.
+This method retrieves the modes one after another by recursively finding the $L$th mode from the residual, assuming the first L1L-1 modes are found after the last turn, until the final residual can be ignored with respect to some criterion set by the algorithm.
+However, since the method is in essence recursive, the accuracy for the next mode to be evaluated is highly depended on the residual determined by previous decomposition in all former turns, which will inevitably introduce accumulated error during the propagation.
+We used official MATLAB implementation on SVMD [25] and observed that for narrow-banded signals wherein the center frequencies of different components are very close, the SVMD method may either introduce some redundant modes or drop some significant modes in their decomposition results, yet the phenomenon is hardly observed in our algorithm.
+Figure 11: The decomposed spectrums for signal $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$ by SVMD under different stopping criteria by considering noise(northwest), exact decomposition(northeast), Bayesian Estimation(southwest), power of the last mode(southeast).
+Figure 12: The original spectrums for signal of the 102th record in MIT-BIH Arrhythmia Database.
+In this section, we pick 2 typical signals for illustration.
+The first one is $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, which we have investigated in the experiments for finding the supporting baseline.
+For the completeness of our comparison, we investigated all stopping criterion implemented in the source code of SVMD.
+The method considering about eliminating noise(northwest) drops the base-frequency and a batch of high frequency, while method for extract reconstruction, method based on Beyesian Estimation, or method considering the power of last mode all result in some redundant modes with low energy on the decomposed spectrum (see colored spectrum at the bottom in each subplot), which are not the essential modes that should included in the decomposition result (Fig.
+11).
+In contrast, our method(Fig.
+10) accurately extracts all necessary modalities without any redundant ones.
+Figure 13: The decomposed spectrums for the V2 lead of the 102th record in MIT-BIH Arrhythmia Database by our method.
+Figure 14: The decomposed spectrums for signal with respect to the V2 lead of the 102th record in MIT-BIH Arrhythmia Database, and decomposed by [25] under different stopping criteria by considering noise(northwest), Bayesian Estimation(southwest), power of the last mode(southeast).
+Note that the figure for the exact decomposition is missed since the SVMD failed to converge in 10 seconds.
+The second one is ECG signals from V2 lead of the 102th record in MIT-BIH Arrhythmia Database [23, 24].
+The experiment result also shows similar phenomenon as stated in the former experiment, the method for exact decomposition fails to converge in more than 10 seconds so we cannot draw the spectrum of the corresponding figure (Fig.
+14).
+Our method in this case can also succeed in retrieving all the effective modes(Fig.
+13).
+4.4 Experiment on Real Signals
+Figure 15: The source signal (blue line) and the reconstructed signal (orange line) from the sum of IMFs obtained from our algorithm with respect to MLII lead from the 100th record.
+Figure 16: The source signal (blue line) and the reconstructed signal (orange line) from the sum of IMFs obtained from our algorithm with respect to V4 lead from the 124th record.
+In this section we investigate the performance of our method on real signals.
+As ECG signals are typical ones that combined with several narrow-banded components, we choose ECG signals as real signals for verification.
+Our ECG signals derives from MIT-BIH Arrhythmia Database [23, 24], and we choose randomly two signals for analysis, wherein the first is the MLII lead from the 100th record and the second is the V4 lead from the 124th record for our experiments respectively.
+Since each original signal has about 65000 samples which makes redundant for our experiment, we choose the first 2000 samples and do down-sampling at the ratio 1:10 for computation efficiency.
+For the limitation of the page we put the course of iteration and the decomposition result in Supplementary Materials(Fig.
+S3.16, Fig.
+S3.17 and Fig.
+S3.18, Fig.
+S3.19 for iteration and decomposition result with respect to the first and second chosen signals respectively), and instead we analyse the correlation coefficient(CC, the higher the better) and the ratio between the power ratio(PR) with respect to the residual and the original signal(the lower the better), and illustrate the gap between the reconstructed signal from the sum of the IMFs that acquired by our algorithm and the source signal.
+For real signals composed of narrow-banded components, our algorithm can obtain effective components, since on one side the gap between the source signal(solid line in Fig.
+15 and Fig.
+16) and reconstructed signal(dashed line in Fig.
+15 and Fig.
+16) is typically small, meaning that one can always revive the source signal with small error by adding all the IMFs obtained from algorithm.
+On the other side, as shown in Table 2, the sum of IMF components obtained by our algorithm can achieve high correlation coefficient(both over 0.99) and low power ratio with respect to residual and the source signal(at the order of $10^{-4}$), which indicates our algorithm can automatically retrieve almost all the meaningful components to the source signal and seldom ignore the significance of components with small energy.
+Table 2: The number of IMFs acquired from our algorithm, the CC(Correlation Coefficients between the reconstructed signal and source signal) and the PR(Power Ratio with respect to the residual and the source signal) result.
+Experiment Number of IMFs $C$ PR
+MLII lead from the 100th record 15 0.9989 4.2635e-04
+V4 lead from the 124th record 14 0.9959 2.0626e-04
 5 Conclusion
@@
 This study is supported by the National Natural Science Foundation of China under Grant No. 61773290 and the Fundamental Research Funds for the Central Universities (22120230311).
-SUPPLEMENTARY MATERIAL
-1 Introduction of VMD
-The key idea of VMD is to decompose an input signal into a series number of sub-signals via variational optimization method in spectral domain[5].
-It is assumed the frequency of each mode is almost compact around a center pulsation and the associated analytic signal is computed by means of Hilbert Transform to get the unilateral frequency spectrum.
-Then each mode’s frequency spectrum was shifted into baseband by multiplying a frequency modulation factor with respect to the estimated center frequency, transforming each band-pass mode into low pass, and the bandwidth can be estimated by $H^{1}$ Gaussian smoothness.
-Afterwards, they decompose the original signal into each modes by minimizing the sum of the bandwidth of each mode with the constraints that the summation of all modes is the original signal.
-In short, their key variation optimization can be listed as following:
-$\min\limits_{u_{k},\omega_{k}}\left\{\sum\limits_{k}\left\|\partial_{t}\left[\left(\delta(t)+\frac{\text{j}}{\pi t}\right)*u_{k}(t)\right]e^{-\text{j}\omega_{k}t}\right\|_{2}^{2}\right\}$
-$\text{s.t.}\quad\sum\limits_{k}u_{k}=f$
-so that each mode can be altered into an associated analytical signal by using Hilbert Transform and shifted into base-band for the following estimation of bandwidth using H1 Gaussian smoothness with the shifted demodulated components.
-To solve the problem, one can construct an augmented Lagrangian Function by introducing both quadratic penalty and Lagrangian Multipliers as follows:
-$\mathcal{L}(u_{k},\omega_{k},\lambda)$ $=\alpha\sum_{k}\left\|\partial_{t}\left[\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)\right]e^{-j\omega_{k}t}\right\|_{2}^{2}$
-$\quad\quad\quad\quad\quad\quad\quad+\left\|f-\sum u_{k}\right\|_{2}^{2}+\langle\lambda,f-\sum u_{k}\rangle$
-This problem can be solved by using ADMM(Alternate Direction Method of Multipliers), in which each component can be updated by solving the equivalent minimization problem:
-$u_{k}^{n+1}=\underset{u_{k}\in\mathbb{R}}{\arg\min}\left\{\alpha\left\|\partial_{t}\left[\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)\right]e^{-j\omega_{k}t}\right\|_{2}^{2}+\left\|f-\sum u_{i}+\frac{\lambda}{2}\right\|_{2}^{2}\right\}.$
-One can rewrite it in frequency domain form, so that convolution becomes multiplication for the convenience of the calculation in subsequent variational operations:
-$\hat{u}_{k}^{n+1}=\underset{\hat{u}_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\alpha\left\|j\omega\left[1+\text{sgn}(\omega+\omega_{k})\right]\hat{u}_{k}(\omega+\omega_{k})\right\|_{2}^{2}+\left\|\hat{f}-\sum\hat{u}_{i}+\frac{\lambda}{2}\right\|_{2}^{2}\right\}.$
-and by substitution of $\omega\rightarrow\omega+\omega_{k}$ to perform a translation and considering the Hermitian symmetry of spectrum with respect to the real signals, both terms can be transformed into half-space integrals with non-negative frequencies:
-$\hat{u}_{k}^{n+1}=\underset{u_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\int_{0}^{\infty}4\alpha(\omega-\omega_{k})^{2}|\hat{u}_{k}(\omega)|^{2}d\omega+2\int\left(\hat{f}-\sum\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)^{2}d\omega\right\}$
-which can be evaluated by vanishing the first variation of positive frequencies for both signal components and center frequencies
-$\hat{u}_{k}^{n+1}=\left(\hat{f}-\sum_{i\neq k}\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)\frac{1}{1+2\alpha(\omega-\omega_{k})^{2}}$
-$\omega_{k}^{n+1}=\frac{\displaystyle\int_{0}^{\infty}\omega\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}{\displaystyle\int_{0}^{\infty}\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}$
-2 Finite Discrete Expression and Their Matrices
+Appendix A Finite Discrete Expression and Their Matrices
 Here we show finite discrete expression and the corresponding the conversion matrices with respect to (19),(20).
@@
 (S1.4) $g^{(4)}(x)\approx\cfrac{g(x+2h)-4g(x+h)+6g(x)-4g(x-h)+g(x-2h)}{h^{4}}$
-(S2.5) $\boldsymbol{G}^{(1)}=\cfrac{1}{12h}\begin{bmatrix}1&-8&0&8&-1&0&\cdots&\cdots&\cdots&0\\ 0&1&-8&0&8&-1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&1&-8&0&8&-1&0\\ 0&0&\cdots&\cdots&0&1&-8&0&8&-1\\ 12h&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&12h\\ -12&12&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-12&12\end{bmatrix}$
-(S2.6) $\boldsymbol{G}^{(2)}=\cfrac{1}{12h^{2}}\begin{bmatrix}-1&16&-30&16&-1&0&\cdots&\cdots&\cdots&0\\ 0&-1&16&-30&16&-1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&-1&16&-30&16&-1&0\\ 0&0&\cdots&\cdots&0&-1&16&-30&16&-1\\ 12h^{2}&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&12h^{2}\\ -12h&12h&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-12h&12h\end{bmatrix}$
-(S2.7) $\boldsymbol{G}^{(3)}=\cfrac{1}{2h^{3}}\begin{bmatrix}-1&2&0&-2&1&0&\cdots&\cdots&\cdots&0\\ 0&-1&2&0&-2&1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&-1&2&0&-2&1&0\\ 0&0&\cdots&\cdots&0&-1&2&0&-2&1\\ 2h^{3}&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&2h^{3}\\ -2h^{2}&2h^{2}&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-2h^{2}&2h^{2}\end{bmatrix}$
-(S2.8) $\boldsymbol{G}^{(4)}=\cfrac{1}{h^{4}}\begin{bmatrix}1&-4&6&-4&1&0&\cdots&\cdots&\cdots&0\\ 0&1&-4&6&-4&1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&1&-4&6&-4&1&0\\ 0&0&\cdots&\cdots&0&1&-4&6&-4&1\\ h^{4}&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&h^{4}\\ -h^{3}&h^{3}&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-h^{3}&h^{3}\end{bmatrix}$
-where the last four lines in matrices correspond to the boundary condition and the conversion matrices $\boldsymbol{A}^{(n)}$ have the same structure.
-3 About Component-Wise Multiplication
-Our component-wise multiplication is not necessarily the Hadamard production of matrices, in that we add some special rules in dealing with column vector, to say in detail, $\forall\boldsymbol{A}_{m\times n},\boldsymbol{v}_{m\times 1}$, we have
-(S3.1) $\boldsymbol{v}\odot\boldsymbol{A}=\boldsymbol{A}\odot\boldsymbol{v}=\boldsymbol{A}*\boldsymbol{V}_{m\times n}$
-where * stands for standard Hadamard production and each column of $\boldsymbol{V}$ is just a copy of $\boldsymbol{v}$. In this way, we can transform (13) into
-(S3.2) $((2\boldsymbol{A^{(2)}\alpha})\odot\boldsymbol{G^{(2)}})g+((4\boldsymbol{A^{(1)}\alpha})\odot\boldsymbol{G^{(3)}})g+((2\boldsymbol{\alpha})\odot\boldsymbol{G^{(4)}})g=\boldsymbol{\beta-\lambda+\mu}$
-By applying this rule we have $(\boldsymbol{A}\boldsymbol{a})\odot(\boldsymbol{B}\boldsymbol{b})=((\boldsymbol{A}\boldsymbol{a})\odot\boldsymbol{B})\boldsymbol{b}$ so that we can treat (13) as common matrix equation and get the solution.
-4 Supplementary Material of the Proof
-4.1 Proof of Convexity
-Let $\mathcal{L}(g,\lambda,\mu,\alpha,\beta)=J_{1}(g(x))+J_{2}(g(x))$, where $J_{1}(g(x))=\displaystyle\int_{\Omega}\alpha(x)g^{\prime\prime 2}(x)\text{d}x$ and $J_{2}(g(x))=\displaystyle-\int_{\Omega}\beta(x)g(x)\text{d}x$.
-For $J_{1}(g(x))$, we have
-(S4.1) $\lambda J_{1}(g_{1}(x))+(1-\lambda)J_{1}(g_{2}(x))-J_{1}[\lambda g_{1}(x)+(1-\lambda)g_{2}(x)]$
+(S1.5) $\boldsymbol{A}^{(1)}=\boldsymbol{G}^{(1)}=\cfrac{1}{12}\begin{bmatrix}1&-8&0&8&-1&0&\cdots&\cdots&\cdots&0\\ 0&1&-8&0&8&-1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&1&-8&0&8&-1&0\\ 0&0&\cdots&\cdots&0&1&-8&0&8&-1\end{bmatrix}$
+(S1.6) $\boldsymbol{A}^{(2)}=\boldsymbol{G}^{(2)}=\cfrac{1}{12}\begin{bmatrix}-1&16&-30&16&-1&0&\cdots&\cdots&\cdots&0\\ 0&-1&16&-30&16&-1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&-1&16&-30&16&-1&0\\ 0&0&\cdots&\cdots&0&-1&16&-30&16&-1\end{bmatrix}$
+(S1.7) $\boldsymbol{A}^{(3)}=\boldsymbol{G}^{(3)}=\cfrac{1}{2}\begin{bmatrix}-1&2&0&-2&1&0&\cdots&\cdots&\cdots&0\\ 0&-1&2&0&-2&1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&-1&2&0&-2&1&0\\ 0&0&\cdots&\cdots&0&-1&2&0&-2&1\end{bmatrix}$
+(S1.8) $\boldsymbol{A}^{(4)}=\boldsymbol{G}^{(4)}=\begin{bmatrix}1&-4&6&-4&1&0&\cdots&\cdots&\cdots&0\\ 0&1&-4&6&-4&1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&1&-4&6&-4&1&0\\ 0&0&\cdots&\cdots&0&1&-4&6&-4&1\end{bmatrix}$
+and after evaluating $(2\boldsymbol{A}^{(2)}\boldsymbol{\alpha})\odot\boldsymbol{G}^{(2)}+(4\boldsymbol{A}^{(1)}\boldsymbol{\alpha})\odot\boldsymbol{G}^{(3)}+2\boldsymbol{\alpha}\odot\boldsymbol{G}^{(4)}$, we introduce another 4 lines depicting the boundary value condition, noted as $\boldsymbol{B}$:
+(S1.9) $\boldsymbol{B}=\begin{bmatrix}1&0&0&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&0&0&1\\ -1&1&0&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&0&-1&1\end{bmatrix}$
+and (19) is then transformed into
+(S1.10) $\begin{bmatrix}(2\boldsymbol{A}^{(2)}\boldsymbol{\alpha})\odot\boldsymbol{G}^{(2)}+(4\boldsymbol{A}^{(1)}\boldsymbol{\alpha})\odot\boldsymbol{G}^{(3)}+2\boldsymbol{\alpha}\odot\boldsymbol{G}^{(4)}\\ \boldsymbol{B}\end{bmatrix}\boldsymbol{g}=\boldsymbol{c}$
+in which
+(S1.11) $\boldsymbol{\alpha}=\begin{bmatrix}\alpha_{0}\\ \alpha_{1}\\ \vdots\\ \alpha_{N-1}\\ \alpha_{N}\end{bmatrix},\boldsymbol{g}=\begin{bmatrix}g_{0}\\ g_{1}\\ \vdots\\ g_{N-1}\\ g_{N}\end{bmatrix}\text{and}\ \ \boldsymbol{c}=\begin{bmatrix}h^{4}c_{2}\\ h^{4}c_{3}\\ \vdots\\ h^{4}c_{N-3}\\ h^{4}c_{N-2}\\ c_{0}\\ c_{N}\\ c_{1}-c_{0}\\ c_{N}-c_{N-1}\end{bmatrix}$
+and $c_{i}=\beta_{i}-\lambda_{i}+\mu_{i}$. In this way the condition number of the matrix to be inverted during evaluation is more gentle.
+Appendix B Supplementary Material of the Proof
+B.1 Proof of Convexity
+Let $\mathcal{O}(g(x))=\mathcal{O}_{1}(g(x))+\mathcal{O}_{2}(g(x))$, where
+$\mathcal{O}_{1}(g(x))=\displaystyle\int_{\Omega}\alpha(x)g^{\prime\prime 2}(x)\text{d}x$ and $\mathcal{O}_{2}(g(x))=\displaystyle-\int_{\Omega}\beta(x)g(x)\text{d}x$.
+For $\mathcal{O}_{1}(g(x))$, we have
+(S2.1) $\lambda\mathcal{O}_{1}(g_{1}(x))+(1-\lambda)J_{1}(g_{2}(x))-\mathcal{O}_{1}[\lambda g_{1}(x)+(1-\lambda)g_{2}(x)]$
 $=\lambda(1-\lambda)\int_{\Omega}\alpha(x)[g_{1}^{\prime\prime}(x)-g_{2}^{\prime\prime}(x)]^{2}\text{d}x\geq 0$
-and $J_{2}(g(x))$ is obviously a linear funtional, so we have
-(S4.2) $J[\lambda g_{1}(x)+(1-\lambda)g_{2}(x)]\leq\lambda J[g_{1}(x)]+(1-\lambda)J[g_{2}(x)]$
+and $\mathcal{O}_{2}(g(x))$ is obviously a linear funtional, so we have
+(S2.2) $\mathcal{O}[\lambda g_{1}(x)+(1-\lambda)g_{2}(x)]\leq\lambda\mathcal{O}[g_{1}(x)]+(1-\lambda)\mathcal{O}[g_{2}(x)]$
 so that our objective functional is convex.
-4.2 Proof of the Zero Duality Gap
+B.2 Concavity of Dual Functional $\mathcal{G}$
+In this part we give a short proof to the concavity of $\mathcal{G}$. According to the definition of $\mathcal{G}$, we have
+(S2.3) $\begin{split}&\ \ \ \ \mathcal{G}(\theta\lambda_{1}(x)+(1-\theta)\lambda_{2}(x),\theta\mu_{1}(x)+(1-\theta)\mu_{2}(x))\\ &=\inf\limits_{g(x)\in\mathcal{D}}\mathcal{L}[g(x),\theta\lambda_{1}(x)+(1-\theta)\lambda_{2}(x),\theta\mu_{1}(x)+(1-\theta)\mu_{2}(x)]\\ &=\inf\limits_{g(x)\in\mathcal{D}}\mathcal{L}[\theta g(x)+(1-\theta)g(x),\theta\lambda_{1}(x)+(1-\theta)\lambda_{2}(x),\theta\mu_{1}(x)+(1-\theta)\mu_{2}(x)]\\ &=\inf\limits_{g(x)\in\mathcal{D}}\{\theta\mathcal{L}[g(x),\lambda_{1}(x),\mu_{1}(x)]+(1-\theta)\mathcal{L}[(g(x),\lambda_{2}(x),\mu_{2}(x)]\}\\ &\geq\theta\inf\limits_{g(x)\in\mathcal{D}}\mathcal{L}[g(x),\lambda_{1}(x),\mu_{1}(x)]+(1-\theta)\inf\limits_{g(x)\in\mathcal{D}}\mathcal{L}[g(x),\lambda_{2}(x),\mu_{2}(x)]\\ &=\theta\mathcal{G}(\lambda_{1}(x),\mu_{1}(x))+(1-\theta)\mathcal{G}(\lambda_{2}(x),\mu_{2}(x))\end{split}$
+which concludes the concavity of $\mathcal{G}$.
+B.3 Proof of the Zero Duality Gap
 Now we begin to prove that the duality gap is zero. Let
@@
 then on one side, we must have
-(S4.6) $\mathcal{G}(\lambda^{*}(x),\mu^{*}(x)=\inf\limits_{g(x)}\mathcal{L}(g(x),\lambda^{*}(x),\mu^{*}(x))\leq\inf\limits_{g(x)}\mathcal{O}(g(x))=\mathcal{O}(g^{*}(x))$
+(S2.7) $\mathcal{G}(\lambda^{*}(x),\mu^{*}(x))=\inf\limits_{g(x)}\mathcal{L}(g(x),\lambda^{*}(x),\mu^{*}(x))\leq\inf\limits_{g(x)}\mathcal{O}(g(x))=\mathcal{O}(g^{*}(x))$
 where * stands for the optimum. On the other hand, one can construct general sets
@@
 hold for all $(p(x),q(x),o)$ in $\mathcal{S}$.
-Note that at least $\forall p(x),q(x)\geq 0$, $\mathcal{O}(g^{*}(x))\geq o$, so that the right end in this case is always non-negative.
-To summarize all above, we must have $\lambda(x)\geq 0$ and $\mu(x)\geq 0$, which indicates we can find a solution in dual space making the dual-gap vanish.
+Note that for $p(x),q(x)\geq 0$, there exist points in $\mathcal{S}$ that makes $o\leq\mathcal{O}(g^{*}(x))$, so there exists points that makes the right side of (S2.12) non-negative, this forces $\lambda(x)\geq 0$ and $\mu(x)\geq 0$, which indicates we can find a solution in dual space making the dual-gap vanish.
 Next we prove the convexity of our feasible domain.
 The feasible set is $0\leq g(x)\leq f(x)$, so that $\forall g_{1}(x),g_{2}(x)\in[0,f(x)]$, we have $\lambda g_{1}(x)+(1-\lambda)g_{2}(x)\in[0,f(x)],\forall\lambda\in[0,1]$, which means the feasible set is also convex.
-5 Experiment Detail
-In this section we begin to provide some details of our experiments as complementary.
 Figure S2.1: The iteration for finding the supporting baseline of FFT with respect to time series $y(t)=100\sin 20\pi t$.
 The 1000th(a), 2000th(b), 3000th(c), 4000th(d), 5000th(e) iteration are respectively shown.
+Appendix C Experiment Detail
+In this section we begin to provide some details of our experiments as complementary.
 Figure S3.2: The gap-elimination course for the found spectrum supporting line with series $y(t)=100\sin 20\pi t$.
@@
 (b) is the result after our KDE process to eliminate the gap at the bottom .
-Figure S5.12: (a) and (b)
-Figure S5.13: The decomposition result (a) of $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$ from the subsequent VMD, using the found number of modes and center frequency as initialization, compared with original VMD results with IMFS number 2 (b) and 3 (c) respectively.
-The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are three) and the residual.
-As compared with (b) and (c), for original VMD, both the results has reconstruction error twice as large as ours.
+Figure S3.12: The decomposition result of $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$ from the subsequent VMD, using the found number of modes and center frequency as initialization.
 Figure S3.13: The iteration for finding the baseline of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The 500th(a), 1000th(b), 1500th(c) iterations are respectively shown.
@@
 The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are three) and the residual.
+Figure S3.16: Iterations for finding the supporting baseline without extending the spectrum to MLII lead from record 100 in MIT-BIH Arrhythmia Database.
+The 1000th(a), 2000th(b), 3000th(c) and 4000th(d) iterations are shown respectively.
+The solid line are the original spectrum and the dotted line are the evaluated supporting baseline during the iterations.
+It takes 4140 iterations to achieve the final convergence.
+Figure S3.17: The decomposition result of the MLII lead from record 100 in MIT-BIH Arrhythmia Database.
+The first line is the original signal and the last the residual, with the 2nd to the 16th lines are the decomposed components acquired by our workflow.
+Figure S3.18: Iterations for finding the supporting baseline without extending the spectrum to V4 lead from record 124 in MIT-BIH Arrhythmia Database.
+The 1500th(a), 3000th(b), 4500th(c) and 6000th(d) iterations are shown respectively.
+The solid line are the original spectrum and the dotted line are the evaluated supporting baseline during the iterations.
+It takes 6899 iterations to achieve the final convergence.
+Figure S3.19: The decomposition result of the V4 lead from record 124 in MIT-BIH Arrhythmia Database.
+The first line is the original signal and the last the residual, with the 2nd to the 15th lines are the decomposed components acquired by our workflow.

### v3

@@
 Abstract
-Automatically determining the number of intrinsic mode functions (IMFs) and their center frequencies in Variational Mode Decomposition (VMD) still remains an open mathematical challenge, which often relies on heuristic prior settings, trial-and-error strategies or complex-field based methods that lack of theoretical guarantees on the convergence.
-In this article, we proposed a novel variational method based on real field, focusing on evaluating the supporting baseline of a signal’s spectrum, to further retrieve the intrinsic mode functions.
-Our method treats automatic extraction of modes as a constrained optimization problem that adversarially maximizes the baseline integral while penalizing its curvature, and transform the problem into iteratively solving a fourth-order boundary value problem via Lagrangian duality.
-Furthermore, we establish the rigorous mathematical proof on the global convergence to our algorithm based on dual ascent in function space.
+Automatically determining the number of intrinsic mode functions (IMFs) and their center frequencies in Variational Mode Decomposition (VMD) remains an open mathematical challenge.
+Existing methods often rely on heuristic prior settings, trial-and-error strategies, or recursive extraction procedures that are computationally inefficient and prone to error accumulation, and generally lack theoretical guarantees on convergence.
+In this article, we propose a novel variational framework that endogenously determines the number of modes.
+We observe that any curve lying below the spectral amplitude naturally divides the spectral domain into connected regions where the spectrum rises above the curve, the number of which defines the modal number K[g]K[g]—a topological functional induced by the cutting curve.
+Since $K[g]$ is discontinuous and intractable for direct optimization, we instead seek the optimal cutting curve itself as a continuous variational surrogate: the optimal curve best separates distinct spectral peaks into individual connected regions above it, while merging noise-induced fragments below.
+This surrogate adversarially maximizes the integral of $g$ (encouraging it to rise and support significant peaks) while penalizing its curvature (suppressing excessive undulations that would fragment the spectrum), transforming the problem into iteratively solving a fourth-order boundary value problem via Lagrangian duality.
+We establish a rigorous proof of global convergence for the dual ascent algorithm in function space.
 Comprehensive numerical experiments on artificial and real-world signals including electrocardiogram (ECG) data show that our method can provide accurate estimates of IMFs and center frequencies, and comparison with methods like Successive VMD also shows our advantages in avoiding redundant modes while ensuring the recovery of necessary components, indicating that we have provided a robust, theoretically grounded initialization routine for VMD.
 keywords:
-narrow-banded signals;variational optimization;Intrinsic Mode Function
-{MSCcodes}
-65K10, 65L10, 68U01, 46N10, 34B05
+Variational Optimization;Convex Optimization;Intrinsic Mode Function;Cutting Curve;Variational Mode Decomposition;Signal Processing
 1 Introduction
@@
 and by substitution of $\omega\rightarrow\omega+\omega_{k}$ to perform a translation and considering the Hermitian symmetry of spectrum with respect to the real signals, both terms can be transformed into half-space integrals with non-negative frequencies:
-(6) $\hat{u}_{k}^{n+1}=\underset{u_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\int_{0}^{\infty}4\alpha(\omega-\omega_{k})^{2}|\hat{u}_{k}(\omega)|^{2}d\omega+2\int\left(\hat{f}-\sum\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)^{2}d\omega\right\}$
+(6) $\hat{u}_{k}^{n+1}=\underset{u_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\int_{0}^{\infty}4\alpha(\omega-\omega_{k})^{2}|\hat{u}_{k}(\omega)|^{2}+2\left(\hat{f}-\sum\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)^{2}\text{d}\omega\right\}$
 which can be evaluated by vanishing the first ordered variation of positive frequencies for both signal components and center frequencies
 (7) $\hat{u}_{k}^{n+1}=\left(\hat{f}-\sum_{i\neq k}\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)\frac{1}{1+2\alpha(\omega-\omega_{k})^{2}}$
-(8) $\omega_{k}^{n+1}=\frac{\displaystyle\int_{0}^{\infty}\omega\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}{\displaystyle\int_{0}^{\infty}\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}$
+(8) $\omega_{k}^{n+1}=\frac{\displaystyle\int_{0}^{\infty}\omega\left|\hat{u}_{k}(\omega)\right|^{2}\text{d}\omega}{\displaystyle\int_{0}^{\infty}\left|\hat{u}_{k}(\omega)\right|^{2}\text{d}\omega}$
 Since its establishment, VMD is always playing an important role in modern signal decomposition in many fields[14, 21, 9, 7, 16, 15, 1].
 Although VMD has provided flexibility in decomposing signals, there are factors that hinder its further applications in that the number of intrinsic mode function must be determined manually as a prior in coordination with penalty factor without a common principle[19, 10, 18, 17].
-In order to alleviate the problem, the subsequent research either needs the knowledge to the range of the number of modes ,[11] retrieves the IMFs recursively ,[12] or highly dependent on a predefined window-size to retrieve the peak and valley of spectra [6].
+In order to alleviate the problem, the subsequent research either needs the knowledge to the range of the number of modes[11], retrieves the IMFs recursively[12], or highly dependent on a predefined window-size to retrieve the peak and valley of spectra [6].
 All these methods either require prior or rely on schemes devoid of convergence guarantees, making the results contains spurious or missed modes.
@@
 To close this gap, we introduce a fundamentally different approach for finding the IMFs.
-Instead of retrieving modes in an trial-and-error way or recursively extracting modes, we focus on the spectrum amplitude and reformulate the problem by introducing supporting baseline, a lower-enveloping curve of a signal’s spectrum which can support the spectrum as tightly as possible while keeping itself as smooth as possible, and transform the problem for finding IMFs into first evaluating the supporting baseline by adversarially maximizing its integral while penalizing its curvature, then retrieving the separated peaks of the original spectrum above the supporting baseline.
+Instead of retrieving modes in an trial-and-error way or recursively extracting modes, we focus on the spectrum amplitude and reformulate the problem by introducing Cutting Curve, a lower-enveloping curve of a signal’s spectrum which can support the spectrum as tightly as possible while keeping itself as smooth as possible, and transform the problem for finding IMFs into first evaluating the Cutting Curve by adversarially maximizing its integral while penalizing its curvature, then retrieving the separated peaks of the original spectrum above the Cutting Curve.
 Our main contribution can be summarized as 4 folds.
-First, we give a new perspective for determining the number of IMFs by introducing supporting baseline of spectrum and solve the problem in real field.
+First, we give a new perspective for determining the number of IMFs by introducing Cutting Curve of spectrum and solve the problem in real field.
 Second, we derive its optimality conditions, establishing a rigorous equivalence between this variational problem and a fourth-order boundary value problem, and establish rigorous mathematical proof for the global convergence to the course of dual ascent iteration in function space.
@@
 2 Methodology and Convergence Proof
-In this section we introduce our method for estimating the number of IMF modes as well as center frequencies of each band.
-We first introduce the background of the problem, and then we propose the concept of Supporting Baseline of a Function, finally we describe our method for finding the baseline and give the convergence proof for our method.
-Figure 1: Illustration to the concept of supporting baseline: In (a)-(c) the solid line is $y=10\cos\pi x-10\cos 5\pi x+x^{2}$.
-The dashed line in (a) is a proper supporting baseline since it actually fits the lower bound of the solid line as well as ignoring high frequent details of the solid line.
-The dashed line in (b) has a gap to the bottom of the solid line so it is not suitable to be a good supporting baseline.
-In contrast, the dashed line in (c) fits too tightly to the lower bound of the solid line, capturing too much detail of the solid line so that it can be reckoned as the lower envelope rather than a proper supporting baseline.
-2.1 Supporting Baseline of a Function
-Without loss of generality we first introduce the Supporting Baseline of a function.
-For any function $f(x)>0$, the supporting baseline is a curve that tightly fits the lower bound of the function while reflects the overall trend of the lower bound.
-Fig.
-1 gives an intuitive illustration about the conception of supporting baseline.
-As illustrated in Fig.
-1, intuitively the supporting baseline should be tightly close to the lower bound of the function, approaching the infimum of functions as close as possible as well as maintain its own smoothness in order to capture the global trend of the lower bound rather than too much details, as capturing too much details will in turn reduce the robustness for separating the modes since there might be some noise added to the signal in either time or frequency domain.
-2.2 Variational Methods for Finding the Supporting Baseline
-To solve the problem for finding the supporting baseline, we use optimization method mainly based on variation and convex optimization[8, 4].
-Let $f(x)\geq 0$ be a known function, the problem of finding the supporting baseline curve of $f(x)$ is equivalent to finding a function $g(x)$ that tightly approaching the point-wise infimum of $f(x)$ while maintaining the global trend of its lower bound.
+According to the theory of Variational Mode Decomposition, each intrinsic mode function should manifest in the spectral domain as a prominent convex bulge, or alternatively as a cluster of adjacent minor bulges.
+In an ideal, flat spectrum, identifying such bulges would be straightforward.
+However, real-world spectra invariably exhibit an underlying trend—a slowly varying floor upon which the modal peaks sit.
+This trend obscures the boundaries between modes, rendering the extraction of bulges far from trivial.
+In this section, we develop a principled framework to address this difficulty.
+Our strategy proceeds as follows.
+First, we formalize the intuitive conditions that an optimal modal partition should satisfy: each detected mode must accumulate sufficient residual energy with respect to the trend of the spectrum; adjacent weak peaks that individually lack saliency should be merged; and distinct modes should be separated by sufficiently deep valleys.
+Second, let $f(x)$ be the spectrum of any investigated signal, we recognize that any curve $g(x)$ lying below $f(x)$ will introduce a partition with respect to the frequency-axis into connected intervals where f(x)g(x)>0f(x)-g(x)>0.
+The number of each connected intervals defines the modal number $K(g(x))$ which is a general function of $g(x)$, and the problem of determining the optimal number $K$ is then recast to selecting an optimal curve $g(x)$ to cut the spectrum, or the optimal Cutting Curve $g^{*}(x)$.
+Since $K(g(x))$ is a discontinuous functional that is hard to be optimized, we construct an surrogate continuous functional which can automatically yields a cutting curve satisfying our assumption that mentioned above.
+The surrogate functional defines a conditional variational problem that adversarially maximizing the integral of the curve and minimizing the curvature.
+By finding the optimal curve defined by the surrogate function satisfying the assumption of optimal $K$, one can evaluate the number of modes in a continuous functional framework, which is much more easier than imposed in the discontinuous circumstances.
+Additionally, we can prove that the optimization of the surrogate functional is a convex problem which can be transformed into an variational problem, and be transformed into a fourth-ordered boundary-value-problem, and a projected dual-ascent algorithm is developed to get the optimal solution.
+We can also establish the global convergence to the unique minimizer in the function space.
+The remainder of the part provides the mathematical detail of the proposed method.
+2.1 Modal Number as a Functional of the Cutting Curve
+In this section we introduce Cutting Curve of a Spectrum.
+Definition 1: Cutting Curve and Support Interval Sets
+Let $f(x)\geq 0$ be any amplitude spectrum of a signal defined on a closed interval $\Omega\subset\mathbb{R}$, for any continuous function $g(x)\in C(\Omega)$ satisfying $0\leq g(x)\leq f(x)$, we define the supporting intervals $\mathcal{S}(g(x))$ as
+(9) $\mathcal{S}(g(x))=\{x\in\Omega|0\leq g(x)\leq f(x)\}$
+and $g(x)$ as a cutting curve, then $\mathcal{S}(g(x))$ can be uniquely decomposed into a series countably many disjoint open sets
+(10) $\mathcal{S}(g(x))=\bigcup_{k=1}^{K(g(x))}C_{k}$
+where
+(11) $C_{i}\cap C_{j}=\varnothing,1\leq i\leq j\leq K(g(x))$
+Then for any cutting function $g(x)$, $K$ is the uniquely functional with respect to $g(x)$ which can be defined as the number of connected components of the set $\mathcal{S}$.
+Definition 2: Optimal Modal Number $K^{*}$
+The best number of modal should satisfy the following conditions:
+(a) Each component $C_{k}$ of $\mathcal{S}(g(x))$ should support sufficient energy of the spectrum, that is the residual $\int_{C_{k}}f(x)-g(x)\text{d}x$ should be sufficiently significant, with respect to the whole trend of the spectrum.
+(b) Between distinct modal components, the cutting curve should remain sufficiently smooth, avoiding spurious oscillations induced by noise or minor fluctuations.
+(c) Adjacent weak peaks lacking individual saliency should not constitute separate modes.
+They are to be either aggregated into a combined mode or, where a dominant peak is present, assimilated as subsidiary ripples of that principal mode.
+As for assumption (a), one should find a curve that approaches the infinitum $f(x)$, yet maximizing the residual $\int f(x)-g(x)\text{d}x$, which in turn encourage the evaluated curve to isolate different modes.
+However, purely encouraging the curve to approach $f(x)$ everywhere will result in nothing to be retrieved, so that the curve should remain at the bottom of $f(x)$ at the interval where peaks exist.
+By considering (b) and (c), the penalty to the curve is a good choice to satisfy the assumption.
+Thus by encouraging the integral of $g(x)$ and discouraging the curvature of $g(x)$, evaluating the optimal $g^{*}(x)$ is implicitly evaluating the best mode number $K$ under our basic assumption.
+However, directly optimization to $K(g(x))$ is difficult since this is not a continuous functional with respect to $g(x)$, so in the following sections we instead optimize $g(x)$ to approach the target.
+Figure 1: Illustration to the concept of best Cutting Curve: In (a)-(c) the solid line is $y=10\cos\pi x-10\cos 5\pi x+x^{2}$.
+The dashed line in (a) is a proper Cutting Curve since it actually fits the lower bound of the solid line as well as ignoring high frequent details of the solid line.
+The dashed line in (b) has a gap to the bottom of the solid line so it is not suitable to be a good cutting curve.
+In contrast, the dashed line in (c) fits too tightly to the lower bound of the solid line, capturing too much detail of the solid line so that it can be reckoned as the lower envelope rather than a proper Cutting Curve.
+2.2 Best Cutting Curve of a Function
+Without loss of generality we first introduce the Cutting Curve of a function.
+For any function $f(x)>0$, the best Cutting Curve is a curve that tightly fits the lower bound of the function while reflects the overall trend of the lower bound.
+Fig.1 gives an intuitive illustration about the conception of Cutting Curve.
+As illustrated in Fig.1, intuitively the Cutting Curve should be tightly close to the lower bound of the function, approaching the infimum of functions as close as possible as well as maintain its own smoothness in order to capture the global trend of the lower bound rather than too much details, as capturing too much details will in turn reduce the robustness for separating the modes since there might be some noise added to the signal in either time or frequency domain.
+2.3 Variational Methods for Finding the Best Cutting Curve
+To solve the problem for finding the best Cutting Curve, we use optimization method mainly based on variation and convex optimization[8, 4].
+Let $f(x)\geq 0$ be a known function, the problem of finding the Cutting Curve curve of $f(x)$ is equivalent to finding a function $g(x)$ that tightly approaching the point-wise infimum of $f(x)$ while maintaining the global trend of its lower bound.
 This can be viewed as the following optimization problem:
@@
 where 0 stands for the left end and $n$ the right end of the curve.
-To solve the boundary value problem of ODE, we use Finite Difference Method.[20] Note that for a differential function $g(x)$, it’s derivative of first to fourth order can be expressed as the linear combination of $g(x)$ itself in the corresponding differential format and we can take similar operation to the function $\alpha(x)$ (See Supplementary Materials A), so that different orders of derivative with respect to $\alpha(x)$ and $g(x)$ can be expressed as the linear combination of themselves, one can easily transform (19) into the matrix-form as follows:
+To solve the boundary value problem of ODE, we use Finite Difference Method[20].
+Note that for a differential function $g(x)$, it’s derivative of first to fourth order can be expressed as the linear combination of $g(x)$ itself in the corresponding differential format and we can take similar operation to the function $\alpha(x)$ (See Supplementary Materials A), so that different orders of derivative with respect to $\alpha(x)$ and $g(x)$ can be expressed as the linear combination of themselves, one can easily transform (22) into the matrix-form as follows:
 (24) $(2\boldsymbol{A^{(2)}\alpha})\odot(\boldsymbol{G^{(2)}g})+(4\boldsymbol{A^{(1)}\alpha})\odot(\boldsymbol{G^{(3)}g})+(2\boldsymbol{\alpha})\odot(\boldsymbol{G^{(4)}g})=\boldsymbol{\beta-\lambda+\mu}$
 […]
-Algorithm 1 The Key Steps for Evaluating Support Baseline
+Algorithm 1 The Key Steps for Evaluating Optimal Cutting Curve
 1: Input $\{f(x)\}$
@@
 In practice, for discrete and potentially noisy spectral data, $f(x)$ can be considered as a smooth interpolation of the observed points.
-Crucially, the core object of interest—the supporting baseline—is intrinsically a smooth function due to the curvature penalty term in the objective.
+Crucially, the core object of interest—the Cutting Curve—is intrinsically a smooth function due to the curvature penalty term in the objective.
 Therefore, the analysis under these smoothness assumptions captures the essential behavior of the algorithm.
@@
 […]
+(40) $\begin{bmatrix}\tilde{\lambda}_{k+1}(x)\\ \tilde{\mu}_{k+1}(x)\end{bmatrix}=\begin{bmatrix}\tilde{\lambda}_{k}(x)\\ \tilde{\mu}_{k}(x)\end{bmatrix}+\eta D\mathcal{G}(\lambda_{k}(x),\mu_{k}(x))$
+as iteration step and rearrange, we have
 […]
@@
 CASE I: If there is only one optima in $\mathcal{G}$, then it is trivial to show that all these sub-sequences converges to the unique optimum by the constraints in (55).
-CASE II: If the optimum of $\mathcal{G}$ forms a convex set, and during the iteration the sequence $(\lambda_{k}(x),\mu_{k}(x))$ is approaching and never fall into the optimal set, then from the middle parts of (51) and (45) we know that the value of $\mathcal{G}$ is approaching the minima and the distance between adjacent term of the sequence $(\lambda_{k}(x),\mu_{k}(x))$ and $(\lambda_{k+1}(x),\mu_{k+1}(x))$ is yielding zero.
-Thus we know in this case the iterating sequence in dual space can always terminate to the point as close as possible to one of the optimum, although the classical convergent behaviour for the sequence in this case might not be guaranteed.
+CASE II: If the optimum of $\mathcal{G}$ forms a convex set, and during the iteration the sequence $(\lambda_{k}(x),\mu_{k}(x))$ is approaching and never fall into the optimal set, then from the middle parts of (55) and (49) we know that the value of $\mathcal{G}$ is approaching the maxima and the distance between adjacent term of the sequence $(\lambda_{k}(x),\mu_{k}(x))$ and $(\lambda_{k+1}(x),\mu_{k+1}(x))$ is yielding zero.
+Thus we know in this case the iterating sequence in dual space can always terminate to the point as close as possible to one of the optimum, and the iterating sequence is of asymptotic regularity.
 CASE III: If the iterating sequence has approached to the unique optima during some steps, then due to (56) the next iteration will still stay at the unique optima.
-CASE IV: If the optimum of $\mathcal{G}$ forms a convex set, and the iterating sequence has approached to one of the optimum $(\lambda^{*}_{1}(x),\mu^{*}_{1}(x))$ early during some iterating step, then since in the convex set there are at least another optima $(\lambda^{*}_{2}(x),\mu^{*}_{2}(x))$, then we have
-(53) $\mathcal{G}(m\lambda^{*}_{1}(x)$ $+(1-m)\lambda^{*}_{2}(x),m\mu^{*}_{1}(x)+(1-m)\mu^{*}_{2}(x))$
-$=\mathcal{G}(\lambda^{*}_{1}(x),\mu^{*}_{1}(x))=\mathcal{G}(\lambda^{*}_{2}(x),\mu^{*}_{2}(x)),m\in[0,1]$
-since the direction of the line connecting the median point and the initial point in the first-order Taylor expansion formula is consistent with $\left[\begin{array}[]{c}\lambda^{*}_{2}(x)-\lambda^{*}_{1}(x)\\ \mu^{*}_{2}(x)-\mu^{*}_{1}(x)\end{array}\right]$, so we have
-(54) $D\mathcal{G}_{(\lambda_{k}(x),\mu_{k}(x))}(\lambda(x),\mu(x))\left[\begin{array}[]{c}\lambda^{*}_{2}(x)-\lambda^{*}_{1}(x)\\ \mu^{*}_{2}(x)-\mu^{*}_{1}(x)\end{array}\right]=0$
-thus the variation direction is always orthogonal to any lines connected between different optima, as a consequence, when projected again during the next iteration, $(\lambda_{k+1}(x),\mu_{k+1}(x))$ must overlap with $(\lambda_{k}(x),\mu_{k}(x))$, so we have $(\lambda_{k+1}(x),\mu_{k+1}(x))=(\lambda_{k}(x),\mu_{k}(x))$.
+CASE IV: If the optimum of $\mathcal{G}$ forms a convex set, and the iterating sequence has approached to one of the optimum $(\lambda^{*}_{1}(x),\mu^{*}_{1}(x))$ early during some iterating step, then in this case the left side of (49) must be zero since under reasonable step size the value of functional $\mathcal{G}$ cannot decrease, so that the middle part of (49) is zero, indicating the iterating sequences will stay still.
 In all cases we show $(\lambda_{k}(x),\mu_{k}(x))$ can always terminate at the point as close as possible to at least one of the optimum and stay still as long as it arrives the optimum during the iteration even if there are more than one optimum in $\mathcal{G}$.
@@
 Proof 2.11.
-As the solution of (10) is continuously dependent on its coefficients, so we have $\lim_{k\rightarrow+\infty}g_{k}(x)=g^{*}(x)$ when $\{(\lambda_{k}(x),\mu_{k}(x))\}$ when in Case I, III and IV, and $g_{k}(x)$ is approaching $g^{*}(x)$ when $g_{k+1}(x)-g_{k}(x)||$ is approaching zero in Case II respectively.
+As the solution of (13) is continuously dependent on its coefficients, so we have $\lim_{k\rightarrow+\infty}g_{k}(x)=g^{*}(x)$ when $\{(\lambda_{k}(x),\mu_{k}(x))\}$ when in Case I, III and IV, and $g_{k+1}(x)-g_{k}(x)||$ is approaching zero in Case II respectively.
 In all cases we can terminate algorithm and guarantee the $\{g_{k}(x)\}$ terminates to one of the optimum as close as possible.
@@
 In this chapter, we focus on the detail about the implementation and numerical scheme about the algorithm for evaluating problem (13).
-Our workflow consists of 4 interconnected stages: (1) We discretize (19) by using Finite Difference Method,[20] transforming it into a linear system.
+Our workflow consists of 4 interconnected stages: (1) We discretize (22) by using Finite Difference Method[20], transforming it into a linear system.
 (2) Crucially, by expressing the derivations of both solution $g(x)$ and the weight function $\alpha(x)$ as linear combination of their nodal values, we recast the entire ODE into a compact matrix form by introducing component-wise multiplication with a mild extension of Hadamard production.
@@
 (3) To further accelerate the convergence and make algorithm effective when handling high or low-passed signals, a smooth extrapolation technique is applied to the input spectrum at the both ends.
-(4) Finally, after obtaining the optimal baseline, a statistically grounded threshold determined via kernel density estimation (KDE), is employed to automatically identify significant spectral peaks, thereby completing the mode extraction.
+(4) Finally, after obtaining the optimal cutting curve, a statistically grounded threshold is determined via kernel density estimation (KDE) on the residual spectrum to filter out background noise.
+The resulting significant peaks are then aggregated using DBSCAN to merge spatially adjacent minor peaks into a single coherent intrinsic mode, thereby completing the mode extraction.
 The following subsections detail each of these components.
@@
 so that we can treat (24) as common matrix equation and get the solution by simply using matrix inversion with modern mathematical software tookits.
-Figure 2: Iterations for finding the supporting baseline without ((a) and (b)) /with ((c) and (d)) smooth extrapolation to the spectrum when iterated to 500 ((a) and (c)) and 1500 ((b) and (d)) steps for the same spectrum.
-The solid line is the original spectrum and the dotted line is the evaluated supporting baseline.
+Figure 2: Iterations for finding the Cutting Curve without ((a) and (b)) /with ((c) and (d)) smooth extrapolation to the spectrum when iterated to 500 ((a) and (c)) and 1500 ((b) and (d)) steps for the same spectrum.
+The solid line is the original spectrum and the dotted line is the evaluated Cutting Curve.
 With smooth extrapolation to the original spectrum, the speed of convergence has significant improvement.
 3.2 Treatment of Boundary Conditions
-The primal problem requires the supporting baseline $g(x)$ confirming with the value of $f(x)$ and $f^{\prime}(x)$ at both ends as boundary conditions, this conducts linear constraints directly onto the first and last elements to the solution vector of $\boldsymbol{g}$.
+The primal problem requires the Cutting Curve $g(x)$ confirming with the value of $f(x)$ and $f^{\prime}(x)$ at both ends as boundary conditions, this conducts linear constraints directly onto the first and last elements to the solution vector of $\boldsymbol{g}$.
 For low and high pass signals of which the spectrum value are high and drop abruptly at either of the ends, this will make $g(x)$ inherits the trend of $f(x)$ at both ends at some interval so that convergence in this case will be very slow with reasonable grid space.
@@
 2 shows the illustrations.
-3.3 Postprocessing: From Baseline to Mode Identification
-Upon the above course when the algorithm converges, we obtain the optimal discrete baseline vector $\boldsymbol{g}^{*}$ above which significant intrinsic modes corresponds to the parts of the original spectrum $\boldsymbol{f}$ that protrude distinctly.
+3.3 Postprocessing: From Cutting Curve to Mode Identification
+Upon the above course when the algorithm converges, we obtain the optimal discrete vector $\boldsymbol{g}^{*}$ above which significant intrinsic modes corresponds to the parts of the original spectrum $\boldsymbol{f}$ that protrude distinctly.
 However, there are some residual background fluctuations at the bottom that hinders one to objectively distinguish those modes.
@@
 so that we can isolate each mode in the function $f(x)-g^{*}(x)-t^{*}$.
-Figure 3: Illustration of the non-consistence of the error between the converged optimal supporting baseline and the original function.
+After that, we use DBSCAN to merge adjacent peaks with low energy at the bottom for $f(x)-g^{*}(x)-t$.
+Figure 3: Illustration of the non-consistence of the error between the converged optimal Cutting Curve and the original function.
 Fig.
@@
 The whole workflow can be summarized as Algorithm 2.
-We first input the original spectrum $f(x)$, then we initialize weight functions $\alpha(x),\beta(x)$, penalty functions $\lambda(x),\mu(x)$ and the supporting baseline function $g(x)$.
+We first input the original spectrum $f(x)$, then we initialize weight functions $\alpha(x),\beta(x)$, penalty functions $\lambda(x),\mu(x)$ and the Cutting Curve function $g(x)$.
 After that we smoothly extend $f(x)$ at both ends and begin Algorithm 1 until it converges.
-Then we do Kernel Density Estimation to figure out the threshold to eliminate the fluttering noise, so as to isolate band components.
+Then we do Kernel Density Estimation to figure out the threshold to eliminate the fluttering noise, and aggregate adjacent minor peaks into single coherent intrinsic mode by DBSCAN.
 Algorithm 2 The Whole Workflow
@@
 9: Subtract noise-gap threshold from $g(x)$ by $g(x)-t_{\text{best}}$.
-10: Get $k$, $\{\omega_{k}\}$ from non-zero intervals.
-11: Begin VMD procedure [5] with $k$ and $\{\omega_{k}\}$
-12: Output $\{u_{k}\}$
-4 Numerical Results
+10: Use DBSCAN to merge spatially adjacent minor peaks into a single coherent intrinsic mode.
+11: Get $k$, $\{\omega_{k}\}$ from non-zero intervals.
+12: Begin VMD procedure [5] with $k$ and $\{\omega_{k}\}$
+13: Output $\{u_{k}\}$
+4 Experiments
 In this section, we show some experiments to validate the applicability of our methods.
@@
 The stopping criterion is set as $\dfrac{\|g_{n}-g_{n-1}\|}{\|g_{n}\|}<10^{-4}$.
-The number of discrete grid points is fixed at $N=200$ for the normalized frequency domain $[0,1]$.
-Detailed descriptions of the test signals and comparative protocols are provided in the corresponding subsections.
-Figure 4: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) with respect to the error for finding the optimum supporting baseline of FFT with respect to time series $y(t)=100\sin 20\pi t$.
+To adaptively determine the effective frequency band, we compute the cumulative energy distribution of the single-sided power spectrum.
+The cut-off frequency $f_{c}$ is defined such that
+(64) $\cfrac{\int_{0}^{f_{c}}S(f)\text{d}f}{\int_{0}^{+\infty}S(f)\text{d}f}=0.95,$
+which is found by scanning from the Nyquist frequency downward until the energy condition is met.
+The resulting $f_{c}$ effectively isolates the dominant spectral components while discarding high-frequency noise tails.
+The truncated spectrum is then down‑sampled to 100 points, and the optimal cutting curve is determined on this down‑sampled spectrum.
+Figure 4: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) with respect to the error for finding the optimum Cutting Curve of FFT with respect to time series $y(t)=100\sin 20\pi t$.
 The iteration lasts for 5161 steps and costs 1.88 seconds.
-4.1 Experiments on Finding the Supporting Baseline
+4.1 Experiments on Finding the Cutting Curve
 We choose five typical signals including single mode signal, multi-mode signal, piecewise continuous signal, signal that are not necessarily narrow-banded, composite signal in which the center frequency of its components are very close.
-We only illustrate the initial and final state of the iteration for the evaluation of the support baseline, as well as the convergent curve with respect to the error in logarithm to indicate the number of iteration steps in the caption, and leave the detail of decomposition result in C.
+We only illustrate the initial and final state of the iteration for the evaluation of the cutting curve, as well as the convergent curve with respect to the error in logarithm to indicate the number of iteration steps in the caption, and leave the detail of decomposition result in Appendix C.
 Experiment 1: The first example is a single mode signal, on the other words, the signal is narrow-banded and only contains one center frequency.
@@
 The normalized center frequency is 0.05 and 0.1, which is relatively very close to each other.
-However, our algorithm can still survive where center frequencies of different modes are very close, since from 6000th iteration and on, the support baseline forms a small bulge at the close frequency points, which just separates the two frequency points.
-Figure 5: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the supporting baseline of FFT with respect to FFT of $y(t)=10\cos 10\pi t+20\sin 20\pi t$.
+However, our algorithm can still survive where center frequencies of different modes are very close, since from 6000th iteration and on, the cutting curve forms a small bulge at the close frequency points, which just separates the two frequency points.
+Figure 5: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the Cutting Curve of FFT with respect to FFT of $y(t)=10\cos 10\pi t+20\sin 20\pi t$.
 The iteration lasts for 5258 steps and costs 1.89 seconds.
@@
 In our cases, it takes about 3026 iterations to achieve the convergence.
-Figure 6: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the supporting baseline of FFT with respect to time series (62).
-Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed supporting baseline precisely captured the trend of the lower bound so that all the modes can be filtered out.
+Figure 6: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the Cutting Curve of FFT with respect to time series (65).
+Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed Cutting Curve precisely captured the trend of the lower bound so that all the modes can be filtered out.
 The iteration lasts for 3026 steps and costs 1.07 seconds.
@@
 On the other side, from the perspective of reconstruction error, it is shown in this example that, results generated by algorithm in [5] with both 2 and 3 modes has reconstruction error as twice as large as ours, which also indicates that our algorithm has better performance in this case.
-Figure 7: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the baseline of function $\frac{1}{1.2+\cos 2\pi t}+\frac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$.
+Figure 7: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the cutting curve of function $\frac{1}{1.2+\cos 2\pi t}+\frac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$.
 The iteration lasts for 3463 steps and costs 1.25 seconds.
+Figure 8: to be continued
 Figure 8: The decomposition result (a) of $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$ from the subsequent VMD, using the found number of modes and center frequency as initialization, compared with original VMD results with IMFS number 2 (b) and 3 (c) respectively.
@@
 The signal has one low-frequency component and 10 alternating current components with different amplitudes, of which the minimum difference between frequency gaps is only 0.04 after normalization and our algorithm can still survive in figuring out all the components.
-Figure 8: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the baseline of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The iteration lasts for 1768 steps and costs 0.64 seconds.
+Figure 9: The initial (a), final convergent state (b) and convergence curve in logarithmic coordinate system (c) for finding the cutting curve of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The iteration lasts for 1768 steps and costs 0.64 seconds.
 4.2 Experiments on Central Frequencies
@@
 We first get the frequency estimation from our algorithm and the result of pure VMD procedure with the same number of IMFs respectively, and then sort them in order, to observe the difference between the results of ours and the corresponding values from VMD.
-Table 1: The evaluated center frequencies
-Experiment Ours Original VMD Max Difference
-1 0.0550 0.0496 0.0054
-2 0.0300, 0.0575 0.0251, 0.0507 0.0068
-3 0.0100, 0.0775 0.1550, 0.2075 0.0008, 0.0530 0.1510, 0.2 0.0245
-4 0.0125, 0.0875 0.2475 6.46e-5, 0.0067 0.0803 0.1682
-5 0.005,0.0775 0.1, 0.125 0.1550, 0.1800 0.2075, 0.2325 0.2575, 0.2825 0.3075 0.000845, 0.0747 0.0999, 0.1250 0.1510, 0.182 0.223, 0.25 0.2760, 0.303 0.339 0.0315
-Table 1 shows the result, where in each experiment the evaluated center frequencies for both ours and the pure VMD are sorted, and the max difference is computed as $\max\limits_{i}|f_{i}-F_{i}|$, where $f$ stands for our estimated frequencies and $F$ pure VMD’s.
-As shown in Table 1, in most cases, our estimation to the center frequencies has low difference with respect to the pure VMD.
-It can be found that, in Experiment 4, the error between estimated center frequencies of ours and pure VMD’s are larger than the average, however, as analysed in Experiment 4, on one side, the original VMD, no matter given IMFs number 2 or 3, fails to achieve reconstruction error as small as ours, on the other side, as shown in Fig.7c even when both started at IMFs number 3, the original VMD seems produce replicated modes at the Gaussian-alike component, with almost the same amplitude and shape, while ours provide relatively independent components when given our initial estimation for the center frequencies, which seems to indicates that our estimation in this case is more reliable.
+Figure 10: The theoretical frequencies (solid line), our estimated frequencies (dashed line) and the VMD-evaluated frequencies (dash-dotted line) after iteration with respect to Experiment 1-5 (a)-(e).
+Fig.
+10 shows the theoretical center frequencies, our estimated frequencies and VMD-evaluated frequencies with respect to the experimental synthetic signals in order, where the solid line stands for the theoretical frequency of each mode, the dashed line stands for our estimation and the dotted-dash line stands for the ones after VMD’s evaluation, and different colors stands for different decomposed modalities.
+It can be seen that our estimated values are very close to the actual modal frequencies, and the accuracy of the estimation is independent of the number of modes to be decomposed, although for VMD procedure there are cases in which the central frequencies of modes with small energy will deviate from the theoretical.
 Figure 11: The original spectrums for signal $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$.
@@
 In this section, we pick 2 typical signals for illustration.
-The first one is $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, which we have investigated in the experiments for finding the supporting baseline.
+The first one is $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, which we have investigated in the experiments for finding the Cutting Curve.
 For the completeness of our comparison, we investigated all stopping criterion implemented in the source code of SVMD.
 The method considering about eliminating noise(northwest) drops the base-frequency and a batch of high frequency, while method for extract reconstruction, method based on Beyesian Estimation, or method considering the power of last mode all result in some redundant modes with low energy on the decomposed spectrum (see colored spectrum at the bottom in each subplot), which are not the essential modes that should included in the decomposition result (Fig.
-11).
+13).
 In contrast, our method(Fig.
-10) accurately extracts all necessary modalities without any redundant ones.
+12) accurately extracts all necessary modalities without any redundant ones.
 Figure 15: The decomposed spectrums for the V2 lead of the 102th record in MIT-BIH Arrhythmia Database by our method.
-Figure 14: The decomposed spectrums for signal with respect to the V2 lead of the 102th record in MIT-BIH Arrhythmia Database, and decomposed by [25] under different stopping criteria by considering noise(northwest), Bayesian Estimation(southwest), power of the last mode(southeast).
-Note that the figure for the exact decomposition is missed since the SVMD failed to converge in 10 seconds.
+Figure 16: Local zoom of the decomposed spectrums for signal with respect to the V2 lead of the 102th record in MIT-BIH Arrhythmia Database, and decomposed by [25] under different stopping criteria by considering Bayesian Estimation(left), power of the last mode(right).
+Note that the figure for the one considering noise and the one with exact decomposition is missed since the SVMD failed to converge in 10 seconds.
 The second one is ECG signals from V2 lead of the 102th record in MIT-BIH Arrhythmia Database [23, 24].
 The experiment result also shows similar phenomenon as stated in the former experiment, the method for exact decomposition fails to converge in more than 10 seconds so we cannot draw the spectrum of the corresponding figure (Fig.
-14).
+16).
 Our method in this case can also succeed in retrieving all the effective modes(Fig.
-13).
+15).
 4.4 Experiment on Real Signals
-Figure 15: The source signal (blue line) and the reconstructed signal (orange line) from the sum of IMFs obtained from our algorithm with respect to MLII lead from the 100th record.
-Figure 16: The source signal (blue line) and the reconstructed signal (orange line) from the sum of IMFs obtained from our algorithm with respect to V4 lead from the 124th record.
-In this section we investigate the performance of our method on real signals.
-As ECG signals are typical ones that combined with several narrow-banded components, we choose ECG signals as real signals for verification.
-Our ECG signals derives from MIT-BIH Arrhythmia Database [23, 24], and we choose randomly two signals for analysis, wherein the first is the MLII lead from the 100th record and the second is the V4 lead from the 124th record for our experiments respectively.
-Since each original signal has about 65000 samples which makes redundant for our experiment, we choose the first 2000 samples and do down-sampling at the ratio 1:10 for computation efficiency.
-For the limitation of the page we put the course of iteration and the decomposition result in Supplementary Materials(Fig.
-S3.16, Fig.
-S3.17 and Fig.
-S3.18, Fig.
-S3.19 for iteration and decomposition result with respect to the first and second chosen signals respectively), and instead we analyse the correlation coefficient(CC, the higher the better) and the ratio between the power ratio(PR) with respect to the residual and the original signal(the lower the better), and illustrate the gap between the reconstructed signal from the sum of the IMFs that acquired by our algorithm and the source signal.
-For real signals composed of narrow-banded components, our algorithm can obtain effective components, since on one side the gap between the source signal(solid line in Fig.
-15 and Fig.
-16) and reconstructed signal(dashed line in Fig.
-15 and Fig.
-16) is typically small, meaning that one can always revive the source signal with small error by adding all the IMFs obtained from algorithm.
-On the other side, as shown in Table 2, the sum of IMF components obtained by our algorithm can achieve high correlation coefficient(both over 0.99) and low power ratio with respect to residual and the source signal(at the order of $10^{-4}$), which indicates our algorithm can automatically retrieve almost all the meaningful components to the source signal and seldom ignore the significance of components with small energy.
-Table 2: The number of IMFs acquired from our algorithm, the CC(Correlation Coefficients between the reconstructed signal and source signal) and the PR(Power Ratio with respect to the residual and the source signal) result.
-Experiment Number of IMFs $C$ PR
-MLII lead from the 100th record 15 0.9989 4.2635e-04
-V4 lead from the 124th record 14 0.9959 2.0626e-04
+To demonstrate the generality for different application domains, we also evaluate its performance on electrocardiogram (ECG) signals.
+ECG signals are representative examples of real-world non-stationary physiological signals composed of multiple characteristic waveforms (such as the P wave, QRS complex, and T wave), each occupying distinct frequency bands.
+Accurately decomposing an ECG signal into its constituent modes is therefore a challenging yet practically meaningful test for any adaptive mode decomposition framework.
+We apply our method to selected records from the MIT-BIH Arrhythmia Database.
+Specifically, we extract 2000-sample segments from the MLII lead of Record 100, the V4 lead of Record 124, and the V2 lead of Record 102.
+For each signal, our method automatically determines the number of intrinsic mode functions and their corresponding initial center frequencies, which are then applied to VMD for the final decomposition.
+Fig.17-Fig.19 shows the decomposition result of MLII lead of Record 100, the V4 lead of Record 124, and the V2 lead of Record 102.
+In each image, the top subfigure represents the original signal, the bottom subfigure represents the residual signal, and the middle graphs represent the decomposed IMFs.
+(b) stands for the spectrums of each decomposed IMF, (c) shows the final converged cutting curve for the spectral amplitude of each signal, and (d) shows the reconstructed signal (orange) with respect to the source signal (blue).
+It is shown that our algorithm determines 4, 4, 2 modes for these three signals in order.
+At the time domain we can see that in each decomposed IMFs in our chosen number of mode can reflect the basic characteristic that is helpful for the disease diagnosis.
+For example, in Fig.17(a) we see the heartbeat rhythm from IMF 1-3 and IMF 4 amplifies the characteristic before and after the peaks.
+The similar phenomenon also happens in Fig.18(a).
+In Fig.19We see that each modality maintains the significant shape features of the original signal.
+The Correlation Coefficients between the reconstructed and source signal is 0.9993, 0.9986, 0.9998 respectively.
+Fig.20 shows the performance comparison between our preferred mode number and center frequencies with respect to randomly chosen mode number from 2 to 10 for each investigated signal.
+The first line is the result for MLII Lead of Record 100, the second line is the result for V4 Lead of Record 124 and the third line is the result for V2 Lead of Record 102.
+(a),(c),(g) stands for the performance of orthogonality (average value of off-diagonal elements), (b),(e),(h) stands for the Correlation Coefficients performance and (c),(f),(i) stands for Power Ratio of Residual performance.
+The red points stands for the performance under our chosen mode number and center frequencies, while the blue lines stands for the performance from randomly choosing mode number from 2 to 10.
+The first 2 lines shows that our algorithm can choose a reasonable mode number protecting over-decomposition, while the third line shows that the V2 Lead of Record 102 seems under-decomposition.
+However, Fig.21-Fig.23 shows that by artificially increasing the mode number it seems increasing some redundant mode or similar mode.
+In Fig.21 and Fig.22 the first 5 modes are similar in that they all reflect the heart beat rhythm, although they are almost orthogonal due to their frequency increasing relationship with each other.
+However, this does not bring any improvement to the interpretability for the help of disease diagnosis.
+Especially, in Fig.23 we see when setting mode number into 3, the decomposed modes can hardly reflect the basic characteristic of the source signal.
+In contrast, the performance of orthogonality in turn improves with the increasing of the mode number as in Fig.21 and Fig.22 the former 5 modes are almost orthogonal due to their frequency increasing relationship with each other.
+This indicates that evaluating the quality of signal decomposition cannot be solely based on performance indicators, but also on the physical meaning of the actual decomposed signal.
+To some extent, our modal determination method is naturally related to the geometric shape of the signal spectrum, and in most cases, the number of modes is determined based on the identifiability of the spectral peaks.
+The resulting modes can be proven valuable in engineering scenarios.
+Figure 17: The Decomposition result for MLII Lead of Record 100.
+(a) The original signal(top), the decomposed IMFs(middle 4), the residual(bottom), (b) The spectrum of each IMF, (c) The final converged cutting line, (d) The source signal and the reconstruction signal.
+Figure 18: The Decomposition result for V4 Lead of Record 124.
+(a) The original signal(top), the decomposed IMFs(middle 4), the residual(bottom), (b) The spectrum of each IMF, (c) The final converged cuttnig curve, (d) The source signal and the reconstruction signal.
+Figure 19: The Decomposition result for V2 Lead of Record 102.
+(a) The original signal(top), the decomposed IMFs(middle 2), the residual(bottom), (b) The spectrum of each IMF, (c) The final converged cutting curve, (d) The source signal and the reconstruction signal.
+Figure 20: The performance of decomposition of our chosen K and center frequencies for each ECG signal compared with random select K from 2 to 10 to proceed VMD.
+Figure 21: The decomposition result on MLII Lead for Record 100 when K is set over our estimation (for example 7).
+Figure 22: The decomposition result on V4 Lead for Record 124 when K is set over our estimation (for example 7).
+Figure 23: The decomposition result on V2 Lead for Record 102 when K is set over our estimation (for example 3).
 5 Conclusion
-In this article, we proposed a globally convergent method that can automatically determine the number of intrinsic mode functions for Variational Mode Decomposition.
-For narrow-banded signals, this method does not need any prior knowledge about the scope of the IMFs number, and provide non-recursively schema to determine IMFs and corresponding central frequencies simultaneously.
-Experiments show that, our method can provide accurate number of IMFs and corresponding center frequencies even if the center frequencies are very close.
-Our implementation is performant in both accuracy and speed.
-Moreover, rigorous proof is given for the convergence of the method, making it theoretically reliable.
+In this article, we address the long-standing open challenge of automatically determining the number of intrinsic mode functions (IMFs) and their corresponding center frequencies for Variational Mode Decomposition (VMD).
+We first establish a geometric insight into this modal initialization problem: the number of valid IMFs is inherently equivalent to the count of disconnected intervals of the spectrum above a critical cutting curve.
+Based on this insight, we formulate the problem within a discrete functional framework defined by the number of isolated intervals with respect to the cutting curve.
+To tackle the intractable non-continuous and non-differentiable nature of this discrete functional optimization, we introduce the cutting curve as a surrogate to transform the original problem into a continuous functional optimization problem.
+On this basis, we propose a globally convergent method for the automatic determination of IMF number and initial center frequencies, with rigorous mathematical proofs provided for the convexity, strong duality and global convergence of the proposed framework.
+For practical engineering implementation, we further develop an efficient and high-precision numerical scheme for the proposed variational framework.
+Specifically, we convert the original optimal variational differential equation into a matrix form, and separate the unknown variables by introducing an extended Hadamard product with compatible broadcasting rules, which reformulates the problem into a standard system of linear equations.
+We also balance the magnitude order of elements in the coefficient matrix to reduce its condition number, which effectively avoids numerical ill-conditioning and enables efficient, high-precision numerical solving even for long-sequence signals.
+We conduct comprehensive experiments on both synthetic narrowband signals and real-world electrocardiogram (ECG) signals.
+Experimental results demonstrate that our method can reliably extract physically meaningful IMFs from complex signals.
+It is validated that the proposed algorithm can select a reasonable number of modes to avoid over-decomposition while guaranteeing high signal reconstruction accuracy.
+Furthermore, the decomposed modal components derived from our method retain meaningful and discriminative features of the target waveform, which is critical for subsequent signal analysis and interpretation tasks.
+This work provides a theoretically rigorous and practically robust initialization paradigm for adaptive VMD, with the potential to be extended to a wide range of signal processing scenarios.
 6 Acknowledgement
-This study is supported by the National Natural Science Foundation of China under Grant No. 61773290 and the Fundamental Research Funds for the Central Universities (22120230311).
+This study is supported by the National Natural Science Foundation of China under Grant No.
+61773290 and the Fundamental Research Funds for the Central Universities (22120230311) and Tongji University Medicine-X Interdisciplinary Research Initiative (Grant No.
+2025-0554-YB-11).
 Appendix A Finite Discrete Expression and Their Matrices
@@
 The feasible set is $0\leq g(x)\leq f(x)$, so that $\forall g_{1}(x),g_{2}(x)\in[0,f(x)]$, we have $\lambda g_{1}(x)+(1-\lambda)g_{2}(x)\in[0,f(x)],\forall\lambda\in[0,1]$, which means the feasible set is also convex.
-Figure S2.1: The iteration for finding the supporting baseline of FFT with respect to time series $y(t)=100\sin 20\pi t$.
+Figure S2.1: The iteration for finding the cutting curve of FFT with respect to time series $y(t)=100\sin 20\pi t$.
 The 1000th(a), 2000th(b), 3000th(c), 4000th(d), 5000th(e) iteration are respectively shown.
@@
 Figure S3.2: The gap-elimination course for the found spectrum supporting line with series $y(t)=100\sin 20\pi t$.
-(a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function.
+(a) is the result of source spectrum subtracting found cutting curve, and one can see that there is some additional noise-gap in the bottom of line subtracted function.
 (b) is the result after our KDE process to eliminate the gap at the bottom .
@@
 The subfigure from top to bottom are respective the source signal, the decomposed mode(only one) and the residual.
-Figure S3.4: The iteration for finding the supporting baseline of FFT with respect to time series $y(t)=10\cos(10\pi t)+20\sin(20\pi t)$.
+Figure S3.4: The iteration for finding the cutting curve of FFT with respect to time series $y(t)=10\cos(10\pi t)+20\sin(20\pi t)$.
 The 1000th(a), 2000th(b), 3000th(c), 4000th(d), 5000th(e) iterations are respectively shown.
 Figure S3.5: The gap-elimination course for the found spectrum supporting line with series $y(t)=10\cos 10\pi t+20\sin 20\pi t$.
-(a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function.
+(a) is the result of source spectrum subtracting found cutting curve, and one can see that there is some additional noise-gap in the bottom of line subtracted function.
 (b) is the result after our KDE process to eliminate the gap at the bottom .
@@
 The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are two) and the residual.
-Figure S3.7: The iteration for finding the supporting baseline of FFT with respect to time series (62).
+Figure S3.7: The iteration for finding the cutting curve of FFT with respect to time series (65).
 The 1000th(a), 2000th(b), 3000th(c) iterations are respectively shown.
-Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed supporting baseline precisely captured the trend of the lower bound so that all the modes can be filtered out.
+Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed cutting curve precisely captured the trend of the lower bound so that all the modes can be filtered out.
 Figure S3.8: The gap-elimination course for the found spectrum supporting line with series (65).
-(a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function.
+(a) is the result of source spectrum subtracting found cutting curve, and one can see that there is some additional noise-gap in the bottom of line subtracted function.
 (b) is the result after our KDE process to eliminate the gap at the bottom.
@@
 The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are four) and the residual.
-Figure S3.10: The iterations for finding the supporting baseline of FFT with respect to time series $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$ are shown.
+Figure S3.10: The iterations for finding the cutting curve of FFT with respect to time series $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$ are shown.
 The 1000th(a), 2000th(b), 3000th(c) iterations are respectively shown.
 Figure S3.11: The noise-elimination course for the found spectrum supporting line with series $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$.
-(a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function.
+(a) is the result of source spectrum subtracting found cutting curve, and one can see that there is some additional noise-gap in the bottom of line subtracted function.
 (b) is the result after our KDE process to eliminate the gap at the bottom .
 Figure S3.12: The decomposition result of $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$ from the subsequent VMD, using the found number of modes and center frequency as initialization.
-Figure S3.13: The iteration for finding the baseline of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The 500th(a), 1000th(b), 1500th(c) iterations are respectively shown.
+Figure S3.13: The iteration for finding the cutting curve of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The 500th(a), 1000th(b), 1500th(c) iterations are respectively shown.
 Figure S3.14: The gap-elimination course for the found spectrum supporting line with series $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$.
-(a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function.
+(a) is the result of source spectrum subtracting found cutting curve, and one can see that there is some additional noise-gap in the bottom of line subtracted function.
 (b) is the result after our KDE process to eliminate the gap at the bottom .
@@
 The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are three) and the residual.
-Figure S3.16: Iterations for finding the supporting baseline without extending the spectrum to MLII lead from record 100 in MIT-BIH Arrhythmia Database.
-The 1000th(a), 2000th(b), 3000th(c) and 4000th(d) iterations are shown respectively.
-The solid line are the original spectrum and the dotted line are the evaluated supporting baseline during the iterations.
-It takes 4140 iterations to achieve the final convergence.
-Figure S3.17: The decomposition result of the MLII lead from record 100 in MIT-BIH Arrhythmia Database.
-The first line is the original signal and the last the residual, with the 2nd to the 16th lines are the decomposed components acquired by our workflow.
-Figure S3.18: Iterations for finding the supporting baseline without extending the spectrum to V4 lead from record 124 in MIT-BIH Arrhythmia Database.
-The 1500th(a), 3000th(b), 4500th(c) and 6000th(d) iterations are shown respectively.
-The solid line are the original spectrum and the dotted line are the evaluated supporting baseline during the iterations.
-It takes 6899 iterations to achieve the final convergence.
-Figure S3.19: The decomposition result of the V4 lead from record 124 in MIT-BIH Arrhythmia Database.
-The first line is the original signal and the last the residual, with the 2nd to the 15th lines are the decomposed components acquired by our workflow.